### v1

Abstract
In this paper, we consider a class of structured nonsmooth fractional minimization, where the first part of the objective is the ratio of a nonnegative nonsmooth nonconvex function to a nonnegative nonsmooth convex function, while the second part is the difference of a smooth nonconvex function and a nonsmooth convex function. This model problem has many important applications, for example, the scale-invariant sparse signal recovery in signal processing. However, the existing methods for fractional programs are not suitable for solving this problem due to its special structure. We first present a novel nonfractional min-max reformulation for the original fractional program and show the connections between their global (local) optimal solutions and stationary points. Based on the reformulation, we propose an alternating maximization proximal descent algorithm and show its subsequential convergence towards a critical point of the original fractional program under a mild assumption. By further assuming the Kurdyka-Łojasiewicz (KL) property of an auxiliary function, we also establish the convergence of the entire solution sequence generated by the proposed algorithm. Finally, some numerical experiments on the scale-invariant sparse signal recovery are conducted to demonstrate the efficiency of the proposed method.
keywords:
nonsmooth fractional program, min-max optimization, proximal algorithm, scale-invariant sparse recovery, Kurdyka-Łojasiewicz property
{AMS}
90C32, 90C26, 90C30, 65K05
1 Introduction
In this paper, we consider the following structured nonsmooth fractional minimization problem
(1) $\min\left\{\frac{f(x)}{g(x)}+h(x):x\in\Omega\cap\mathcal{C}\right\},$
where $f,g:\mathbb{R}^{n}\rightarrow[0,+\infty)$ and $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are proper closed, $\mathcal{C}\subseteq\mathbb{R}^{n}$ is closed and $\mathcal{C}\cap\Omega\neq\emptyset$ with $\Omega:=\left\{x\in\mathbb{R}^{n}:g(x)\neq 0\right\}$. Throughout this paper, we adopt the following blanket assumption for problem (1).
Assumption 1
  1. (i)
    $f$ is locally Lipschitz continuous on $\Omega$.
  2. (ii)
    $g$ is convex on $\mathbb{R}^{n}$.
  3. (iii)
    $h:=h_{1}-h_{2}$, where $h_{1}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is locally Lipschitz differentiable around each $x\in\Omega$ and $h_{2}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is convex on $\mathbb{R}^{n}$.
  4. (iv)
    The proximal operator associated with the sum of $f$ and the indicator function on $\mathcal{C}$ can be evaluated. [^1] The proximal operator associated with the sum of $f$ and the indicator function on $\mathcal{C}$ is defined as $\mathrm{prox}_{f+\iota_{\mathcal{C}}}(x):=\arg\min\{f(z)+\frac{1}{2}\|x-z\|^{2}_{2}:z\in\mathcal{C}\}.$
Problem (1) covers many important optimization models in machine learning and scientific computing. In this work, we focus on its application to the scale-invariant sparse signal recovery [32, 37, 38, 39, 40, 42, 43]. Let $A\in\mathbb{R}^{m\times n}$ be the sensing matrix, $b\in\mathbb{R}^{m}$ be the possibly noisy measurement, $\underline{x}\in\mathbb{R}^{n}$ and $\overline{x}\in\mathbb{R}^{n}$ be the lower bound and upper bound of the underlying signal, respectively. We are particularly interested in the $L_{1}/L_{2}$ robust signal recovery model
(2) $\min\left\{\frac{\|x\|_{1}}{\|x\|_{2}}+\frac{\lambda}{2}\mathrm{dist}^{2}\left(Ax-b,\mathcal{S}_{\mu}\right):x\neq 0,~{}\underline{x}\leq x\leq\overline{x},~{}x\in\mathbb{R}^{n}\right\},$
and the $L_{1}/S_{K}$ robust signal recovery model
(3) $\min\left\{\frac{\|x\|_{1}}{\|x\|_{(K)}}+\frac{\lambda}{2}\mathrm{dist}^{2}(Ax-b,\mathcal{S}_{\mu}):x\neq 0,~{}\underline{x}\leq x\leq\overline{x},~{}x\in\mathbb{R}^{n}\right\},$
where $\|x\|_{(K)}$ denotes the sum of the $K$ largest absolute values of entries in $x$ for the positive integer $K$, $\lambda>0$, $\mathcal{S}_{\mu}:=\{z\in\mathbb{R}^{m}:\|z\|_{0}\leq\mu\}$ with $\mu$ being a nonnegative integer and $\|\cdot\|_{0}$ counting the number of nonzero elements in the vector. A direct computation yields that $\mathrm{dist}^{2}(Ax-b,\mathcal{S}_{\mu})=\|Ax-b\|^{2}_{2}-\|\mathcal{T}_{\mu}(Ax-b)\|^{2}_{2}$ with $\mathcal{T}_{\mu}(z)$ being the projection operator onto $\mathcal{S}_{\mu}$, which retains the $\mu$ largest elements of the vector $z\in\mathbb{R}^{m}$ in absolute value and sets the rest to be zero. Moreover, it is not hard to verify that $\|\mathcal{T}_{\mu}(Ax-b)\|^{2}_{2}$ is convex. Then, it is obvious that both problems (2) and (3) are special examples of problem (1) with $\mathcal{C}$ being the indicator function on $\{x\in\mathbb{R}^{n}:\underline{x}\leq x\leq\overline{x}\}$, $f(x)=\|x\|_{1}$, $g(x)=\|x\|_{2}$ or $\|x\|_{(K)}$, $h_{1}(x)=\frac{\lambda}{2}\|Ax-b\|^{2}_{2}$, and $h_{2}(x)=\frac{\lambda}{2}\|\mathcal{T}_{\mu}(Ax-b)\|_{2}^{2}$.
Now we briefly review two categories of approaches for solving structured nonsmooth fractional optimization problems. The first category of approaches is based on the classical Dinkelbach’s method [14, 17, 19, 35], which is widely used for single-ratio fractional programs. Note that problem (1) can be naively converted into a signal-ratio optimization problem with the numerator $f+hg$, for which the Dinkelbach’s method generate the new iterate $x^{k+1}$ by solving
(4) $\min\left\{f(x)+(hg)(x)-\theta_{k}g(x):x\in\mathcal{C}\right\},$
where $\theta_{k}\in\mathbb{R}$ is updated via $\theta_{k}:=f(x^{k})/g(x^{k})+h(x^{k})$. This iterative scheme is not practical in general since it may be very difficult and expensive to solve an optimization problem of type (4) in each iteration. Recently, proximal algorithms based on the Dinkelbach’s method have been developed for a class of nonsmooth single-ratio fractional programs programs [8, 9, 25, 44]. If we ignore the line-search and extrapolation involved, then applying these algorithms to problem (1) yields the iteration scheme
(5) $x^{k+1}\in\arg\min\left\{f(x)+(hg)(x)-\left\langle\theta_{k}y^{k},x-x^{k}\right\rangle+\frac{\|x-x^{k}\|_{2}^{2}}{2\eta_{k}}:x\in\mathcal{C}\right\},$
or
(6) $x^{k+1}\in\arg\min\left\{f(x)+\left\langle\nabla(hg)(x^{k})-\theta_{k}y^{k},x-x^{k}\right\rangle+\frac{\|x-x^{k}\|_{2}^{2}}{2\eta_{k}}:x\in\mathcal{C}\right\},$
where $y^{k}\in\partial g(x^{k})$, $\eta_{k}>0$ is some stepsize, and $\theta_{k}$ is defined as in (4). It is worth noting that (5) actually requires that the proximal operator associated with the sum of $f+hg$ and indicator function on $\mathcal{C}$ can be evaluated, while (6) assumes the smoothness of $hg$. Unfortunately, the above requirements do not hold in general and consequently the iteration schemes (5) and (6) are not suitable for tackling problem (1). Therefore, we conclude that the Dinkelbach’s method and its existing variants are not applicable to solving problem (1). The second category of approaches are based on the so-called quadratic transform [5, 10, 36]. This transform introduces an additional variable and provides an equivalent nonfractional maximization problem for fractional programs that maximizes the sum of several ratios. Specifically, it can be applied to the alternative problem that is obtained via replacing the minimum in problem (1) by the maximum, that is,
(7) $\max\left\{\frac{f(x)}{g(x)}+h(x):x\in\Omega\cap\mathcal{C}\right\}.$
Consequently, the equivalent nonfractional problem for (7) is formulated as
(8) $\max\left\{2y\sqrt{f(x)}-y^{2}g(x)+h(x):x\in\Omega\cap\mathcal{C},~{}y\in\mathbb{R}\right\}.$
Since for a fixed $x\in\Omega\cap\mathcal{C}$ the optimal objective of (8) is attained at $y=\sqrt{f(x)}/g(x)$, one can easily utilize block coordinate ascent methods for efficiently solving (8). However, the minimization problem (1) is generally independent of the maximization problem (7) due to the non-negativity requirements of their numerators and denominators. Hence, the quadratic transform can not be directly used to construct an equivalent nonfractional minimization problem for (1).
We also note that some algorithms have been specifically designed for solving the scale-variant sparse signal recovery model, which are based on neither the Dinkelbach’s method nor the quadratic transform. The alternating direction method of multipliers (ADMM) has been applied to the noiseless $L_{1}/L_{2}$ signal recovery model and the model (2) with $\mu=0$ via various variable-splitting schemes [32, 39]. However, these ADMM-type algorithms rely on the structure of the $L_{1}/L_{2}$ function and can not be extended to deal with problem (1) which admits a more general setting for $f$ and $g$. For example, they can not be generalized to solve model (3) due to the nonsmoothness of $\|\cdot\|_{(K)}$. Very recently, the gradient descent flow algorithm (GDFA) is proposed for models (2) and (3) with $\mu=0$ [38]. Although the efficiency of GDFA is demonstrated in the numerical experiments of [38], so far its convergence remains unknown to the best of our knowledge.
Motivated by the above analysis, we aim at designing efficient numerical methods with guaranteed convergence for solving problem (1) by exploiting its intrinsic structure. Specifically, the contributions of this paper are summarized as follows. First, we present an novel nonfractional min-max reformulation for problem (1), which is in the form of
[…]
This reformulation is equivalent to (1) in the sense that $x^{\star}\in\mathbb{R}^{n}$ is a global (local) minimizer of problem (1) if and only if $(x^{\star},c_{\star})$ is a global (local) min-max point of problem (9) for some $c_{\star}\in\mathbb{R}$. Moreover, we prove that $x^{\star}$ is a (Fréchet) stationary point of problem (1) if and only if $(x^{\star},c_{\star})$ is a stationary point of problem (9) for some $c_{\star}\in\mathbb{R}$. To the best of our knowledge, problem (9) is the first equivalent min-max reformulation for fractional programs of the type (1). Second, we propose an alternating maximization proximal descent algorithm (AMPDA) for solving the min-max problem (9). Problem (9) falls into the class of nonsmooth nonconvex-concave min-max problems, which has received considerable attention recently; see, for example, [7, 16, 18, 22, 24, 26] and the references therein. However, existing min-max optimization methods for this class are not suitable for solving problem (9) since for fixed maximization variable $c\in\mathbb{R}$ its objective function involves a nonsmooth nonconvex product term $-c^{2}f(x)g(x)$ with respect to the minimization variable $x$. Hence, our development of the AMPDA is nontrivial. In particular, we ingeniously employ the majorization-minimization scheme and the line-search technique to perform a proximal update associated with the sum of $f$ and the indicator function on $\mathcal{C}$ for the variable $x$, where the line-search procedure determines the proximal stepsize and is terminated when some sufficient descent condition is satisfied. We show that any accumulation point of the solution sequence generated by the AMPDA is a critical point of problem (1) under a mild assumption. Lastly, by further assuming that some auxiliary function is a proper closed KL function, we prove the convergence of the entire solution sequence generated by AMPDA. It is worth noting that our sequential convergence analysis for AMPDA does not require any additional assumption on the denominator $g$, while those for nonsmooth fractional optimization methods in existing literature always assume that either the conjugate of $g$ is continuous on its domain or $g$ is locally Lipschitz differentiable on $\Omega$; see, for example, [9, 25].
The rest of this paper is organized as follows. In Section 2, we introduce notation and present some preliminary materials. In Section 3, we study the relationships between problem (1) and its min-max reformulation (9). We propose the AMPDA and establish its subsequential convergence in Section 4. The sequential convergence of the AMPDA is further investigated in Section 5. Finally, we present some numerical results of the AMPDA for models (2) and (3) in Section 6.
2 Notation and preliminaries
[…]
$\iota_{\mathcal{A}}(x)=\begin{cases}0,&\text{if $x\in\mathcal{A}$,}\\ +\infty,&\text{else.}\end{cases}$
An extended-real-value function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be proper if its domain $\mathrm{dom}(\varphi):=\{x\in\mathbb{R}^{n}:\varphi(x)<+\infty\}$ is nonempty. A proper function $\varphi$ is said to be closed if $\varphi$ is lower semi-continuous on $\mathbb{R}^{n}$. A function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to be locally Lipschitz continuous around $x\in\mathrm{dom}(\varphi)$ with a modulus $L_{x}>0$ (resp., relative to $\mathrm{dom}(\varphi)$), if there exists $\delta_{x}>0$ such that $|\varphi(u)-\varphi(v)|_{2}\leq L_{x}\|u-v\|_{2}$ holds, for all $u,v\in\mathcal{B}(x,\delta_{x})$ (resp., for all $u,v\in\mathcal{B}(x,\delta_{x})\cap\mathrm{dom}(\varphi)$). A function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is said to be locally Lipschitz differentiable around $x\in\mathrm{dom}(\varphi)$ with a modulus $\tilde{L}_{x}>0$, if there exists $\delta_{x}>0$, such that $\varphi$ is continuously differentiable on $\mathcal{B}(x,\delta_{x})$ and $\|\nabla\varphi(u)-\nabla\varphi(v)\|_{2}\leq\tilde{L}_{x}\|u-v\|_{2}$ holds for all $u,v\in\mathcal{B}(x,\delta_{x})$.
2.1 Preliminaries on nonsmooth analysis
We first review some preliminaries on subdifferentials of nonconvex functions [15, 29, 34]. For a proper function $\varphi$, the Fréchet subdifferential at $x\in\mathrm{dom}(\varphi)$ is defined by
^$\widehat{\partial}\varphi(x):=\left\{y\in\mathbb{R}^{n}:\mathop{\lim\inf}\limits_{\begin{subarray}{c}z\to x\\ z\neq x\end{subarray}}\;\frac{\varphi(z)-\varphi(x)-\left\langle{y},{z-x}\right\rangle}{\|z-x\|_{2}}\geq 0\right\},$
and limiting subdifferential at $x\in\mathrm{dom}(\varphi)$ is defined by
${\partial}\varphi(x):=\left\{y\in\mathbb{R}^{n}:\exists x^{k}\to x,~{}\varphi(x^{k})\to\varphi(x),~{}y^{k}\to y\text{ with $y^{k}\in\widehat{\partial}\varphi(x^{k})$ for each $k$}\right\}.$
We set ^$\widehat{\partial}\varphi(x)=\partial\varphi(x)=\emptyset$ by convention when $x\notin\mathrm{dom}(\varphi)$, and define $\mathrm{dom}(\partial\varphi):=\{x:\partial\varphi(x)\neq\emptyset\}$. It is clear that ^$\widehat{\partial}(\alpha\varphi)(x)=\alpha\widehat{\partial}\varphi(x)$ holds for all $x\in\mathrm{dom}(\varphi)$ and $\alpha>0$, and ^$\widehat{\partial}\varphi(x)\subseteq{\partial}\varphi(x)$ holds for all $x\in\mathrm{dom}(\varphi)$. It holds that ^$\widehat{\partial}\varphi(x)=\{\nabla\varphi(x)\}$ if $\varphi$ is differentiable at $x$, and $\partial\varphi(x)=\{\nabla\varphi(x)\}$ holds when $\varphi$ is continuously differentiable around $x$ ([34, Exercise 8.8(b)]). For a convex $\varphi$, the Fréchet and limiting subdifferential both consist with the classical subdifferential of a convex function at each $x\in\mathrm{dom}(\varphi)$ ([34, Proposition 8.12]), that is,
^$\widehat{\partial}\varphi(x)=\partial\varphi(x)=\{y\in\mathbb{R}^{n}:\varphi(z)-\varphi(x)-\left\langle{y},{z-x}\right\rangle\geq 0,\text{~{}for all~{}}z\in\mathbb{R}^{n}\}.$
If $\varphi(x)=\varphi_{1}(x_{1})+\varphi_{2}(x_{2})$ for proper closed functions $\varphi_{1}:\mathbb{R}^{n_{1}}\to(-\infty,+\infty]$ and $\varphi_{2}:\mathbb{R}^{n_{2}}\to(-\infty,+\infty]$, where $x\in\mathbb{R}^{n}$ is expressed as $(x_{1},x_{2})$, then it holds that ^$\widehat{\partial}\varphi(x)=\widehat{\partial}\varphi_{1}(x_{1})\times\widehat{\partial}\varphi_{2}(x_{2})$ ([34, Proposition 10.5]). Moreover, for proper functions $\varphi_{1},~{}\varphi_{2}:\mathbb{R}^{n}\to(-\infty,+\infty]$, there holds ^$\widehat{\partial}\varphi_{1}(x)+\widehat{\partial}\varphi_{2}(x)\subseteq\widehat{\partial}(\varphi_{1}+\varphi_{2})(x)$ ([34, Corollary 10.9]), where the equality holds if $\varphi_{1}$ or $\varphi_{2}$ is differentiable at $x$ ([34, Section 8.8 Exercise]). For a function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$, the generalized Fermat’s Rule states that $0\in\widehat{\partial}\varphi(x^{\star})$ holds if $x^{\star}\in\mathrm{dom}(\varphi)$ is a local minimizer of $\varphi$ ([34, Theorem 10.1]).
We finally review some properties of real-valued convex functions. Let $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ be convex. Then $\varphi$ is continuous on $\mathbb{R}^{n}$ [34, Theorem 2.35]. In view of this, one can verify that $\partial\varphi$ is outer semicontinuous, that is, $\bar{y}\in\partial\varphi(\bar{x})$ holds if there exists some $x^{k}\to\bar{x}$ and $y^{k}\to\bar{y}$ with $y^{k}\in\partial\varphi(x^{k})$ for each $k\in\mathbb{N}$. For any bounded $\mathcal{S}\subseteq\mathbb{R}^{n}$, the $\varphi$ is Lipschitz continuous on $\mathcal{S}$, and $\cup_{x\in\mathcal{S}}\partial\varphi(x)$ is nonempty and bounded ([6, Proposition 5.4.2]). The convex conjugate function $\varphi^{*}:\mathbb{R}^{n}\to[-\varphi(0),+\infty]$ is defined at $y\in\mathbb{R}^{n}$ as $\varphi^{*}(y):=\sup\{\left\langle{x},{y}\right\rangle-\varphi(x):x\in\mathbb{R}^{n}\}$. The conjugate $\varphi^{*}$ is a proper closed convex function [33, Theorem 12.2]. Furthermore, for any $x,y\in\mathbb{R}^{n}$, the following three statements are equivalent ([34, Proposition 11.3]): $\left\langle{x},{y}\right\rangle=\varphi(x)+\varphi^{*}(y)\Leftrightarrow y\in\partial\varphi(x)\Leftrightarrow x\in\partial\varphi^{*}(y)$. Therefore, it is clear that the Fenchel-Young Inequality, i.e., $\varphi(x)+\varphi^{*}(y)\geq\left\langle{x},{y}\right\rangle$, holds for all $x,y\in\mathbb{R}^{n}$, and it becomes an equality when $y\in\partial\varphi(x)$. Besides, from the Fenchel-Young Inequality, we know that $y\in\mathrm{dom}(\varphi^{*})$ if $y\in\partial\varphi(x)$ for some $x\in\mathbb{R}^{n}$.
2.2 KL property
We now review the KL property, which has been used extensively in the convergence analysis of various first order methods.
Definition 2.1 (KL property [1, 2]).
A proper function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is said to satisfy the KL property at $x\in\mathrm{dom}(\partial\varphi)$ if there exist $\epsilon\in(0,+\infty]$, $\delta>0$ and a continuous concave function $\phi:[0,\epsilon)\to[0,+\infty)$ such that:
  1. (i)
    $\phi(0)=0$;
  2. (ii)
    $\phi$ is continuously differentiable on $(0,\epsilon)$ with $\phi^{\prime}>0$;
  3. (iii)
    For any $z\in\mathcal{B}(x,\delta)$ satisfying $\varphi(x)<\varphi(z)<\varphi(x)+\epsilon$, it holds that $\phi^{\prime}(\varphi(z)-\varphi(x))\mathrm{~{}dist}(0,\partial\varphi(z))\geq 1$.
A proper function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is called a KL function, if it satisfies the KL property at each point in $\mathrm{dom}(\partial\varphi)$. For connections between the KL property and the well-known error bound theory [27, 28, 31], we refer the interested readers to [11, 23]. The following proposition provides a framework for proving sequential convergence using the KL property.
Proposition 2.2.
[25, Proposition 2.7] Let $\Psi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to(-\infty,+\infty]$ be proper lower semicontinuous. Consider a bounded sequence $\{(u^{k},v^{k})\in\mathbb{R}^{n}\times\mathbb{R}^{m}:k\in\mathbb{N}\}$ satisfying the following three conditions:
  1. (i)
    (Sufficient descent condition). There exists $a>0$, such that
    $\Psi(u^{k+1},v^{k+1})+a\|u^{k+1}-u^{k}\|_{2}^{2}\leq\Psi(u^{k},v^{k}),\text{~{}~{}~{}~{}~{}for all $k\in\mathbb{N}$;}$
  2. (ii)
    (Relative error condition). There exists $b>0$, such that
    $\mathrm{dist}(0,\partial\Psi(u^{k+1},v^{k+1}))\leq b\|u^{k+1}-u^{k}\|_{2},\text{~{}~{}~{}~{}~{}for all $k\in\mathbb{N}$;}$
  3. (iii)
    (Continuity condition). The limit $\Psi_{\infty}:=\lim_{k\to\infty}\Psi(u^{k},v^{k})$ exists and $\Psi\equiv\Psi_{\infty}$ holds on $\Upsilon$, where $\Upsilon$ denotes the set of accumulation points of $\{(u^{k},v^{k}):k\in\mathbb{N}\}$.
If $\Psi$ satisfies the KL property at each point of $\Upsilon$, then we have $\sum_{k=0}^{\infty}\|u^{k+1}-u^{k}\|_{2}<+\infty$, $\lim_{k\to\infty}u^{k}=u^{\star}$ and $0\in\partial\Psi(u^{\star},v^{\star})$ for some $(u^{\star},v^{\star})\in\Upsilon$.
2.3 Min-max problem
Consider the following min-max problem:
(10) $\min_{x\in\mathcal{F}_{X}}\max_{y\in\mathcal{F}_{Y}}~{}\psi(x,y),$
where $\mathcal{F}_{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{F}_{Y}\subseteq\mathbb{R}^{m}$ are nonempty, and the objective $\psi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is continuous. We assume that $\mathop{\arg\max}_{y\in\mathcal{F}_{Y}}\psi(x,y)\neq\emptyset$ for all $x\in\mathcal{F}_{X}$, and we define an envelop function $\varphi:\mathcal{F}_{X}\to\mathbb{R}$ as
$\varphi(x):=\max_{y\in\mathcal{F}_{Y}}~{}\psi(x,y).$
Problem (10) is called nonconvex-concave if for a fixed $x\in\mathcal{F}_{X}$, $\psi(x,\cdot)$ is concave, and for a fixed $y\in\mathcal{F}_{Y}$, $\psi(\cdot,y)$ is not convex. A point $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is said to be a saddle point of problem (10) if $\psi(x^{\star},y)\leq\psi(x^{\star},y^{\star})\leq\psi(x,y^{\star})$ holds for all $(x,y)\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$. The concept of saddle points is usually used to describe the optimality of the convex-concave min-max problems. Unfortunately, it is well-known that such points may not exist for nonconvex-concave min-max problems. Due to this, the notions of global and local min-max points are introduced in [21, Definition 9] and [21, Definition 14], respectively. Here we consider equivalent definitions of these two concepts given in [21, Remark 10 and Lemma 16] as follows.
Definition 2.3 (global min-max point).
[21, Remark 10] We say that $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is a global min-max point of problem (10), if $y^{\star}$ is a global maximizer of $\psi(x^{\star},\cdot)$ and $x^{\star}$ is a global minimizer of $\varphi$, i.e.,
(11) $\begin{cases}\varphi(x^{\star})\leq\varphi(x),&\text{for all $x\in\mathcal{F}_{X}$,}\\ \psi(x^{\star},y^{\star})\geq\psi(x^{\star},y),&\text{for all $y\in\mathcal{F}_{Y}$.}\end{cases}$
Definition 2.4 (local min-max point).
[21, Lemma 16] We say that $(x^{\star},y^{\star})\in\mathcal{F}_{X}\times\mathcal{F}_{Y}$ is a local min-max point of problem (10), if $y^{\star}$ is a local maximizer of $\psi(x^{\star},\cdot)$ relative to $\mathcal{F}_{Y}$, and there exists an $\epsilon_{0}>0$ such that $x^{\star}$ is always a local minimizer of $\varphi_{\epsilon}$ for all $\epsilon\in(0,\epsilon_{0}]$ relative to $\mathcal{F}_{X}$, where the function $\varphi_{\epsilon}$ is defined as $\varphi_{\epsilon}(x):=\max\{\psi(x,y):y\in\mathbb{R}^{m},\|y-y^{\star}\|_{2}\leq\epsilon\}$.
In view of [21, Proposition 11], problem (10) always has a global min-max point if $\mathcal{F}_{X}$ is compact. A saddle point of problem (10) is also a local min-max point. For more discussions on their relationships, we refer the readers to [20, 21].
3 The connections between problem (1) and the min-max reformulation (9)
In this section, we investigate the connections between global and local optimal solutions as well as stationary points of problems (1) and (9). To this end, we first introduce the extended objective $F:\mathbb{R}^{n}\to(-\infty,+\infty]$ for problem (1):
(12) $F(x):=\begin{cases}\frac{f(x)}{g(x)}+h(x),&\text{if $x\in\Omega\cap\mathcal{C}$},\\ +\infty,&\text{else.}\end{cases}$
Obviously, problem (1) can be rewritten as $\min\{F(x):x\in\mathbb{R}^{n}\}$. Moreover, we use $\widetilde{F}$ to denote the extended objective of the min-max problem (9), i.e.,
(13) $\widetilde{F}(x,c):=2cf(x)-c^{2}f(x)g(x)+h(x)+\iota_{\Omega\cap\mathcal{C}}(x).$
By a direct computation, we obtain that for all $(x,c)\in(\Omega\cap\mathcal{C})\times\mathbb{R}$,
(14) $F(x)-\widetilde{F}(x,c)=f(x)g(x)\left(\frac{1}{g(x)}-c\right)^{2}.$
In view of (14) as well as the non-negativity of $f$ and $g$, we know that for all $x\in\Omega\cap\mathcal{C}$,
(15) $F(x)=\max_{c\in\mathbb{R}}\widetilde{F}(x,c),$
where the maximum is attained at some $c_{x}\in\mathbb{R}$ satisfying $f(x)c_{x}=f(x)/g(x)$. The following proposition shows the relationship between the global minimizers of problem (1) and the global min-max points of problem (9).
Proposition 3.1.
If $x^{\star}$ is a global minimizer of problem (1), then $(x^{\star},1/g(x^{\star}))$ is a global min-max point of problem (9). Conversely, if $(x^{\star},c_{\star})$ is a global min-max point of problem (9), then $x^{\star}$ is a global minimizer of problem (1).
Proof 3.2.
If $x^{\star}$ is a global minimizer of (1), then we have $x^{\star}\in\arg\min\{F(x):x\in\Omega\cap\mathcal{C}\}$. This, together with (15), leads to
(16) $x^{\star}\in\arg\min\left\{\max_{c\in\mathbb{R}}\widetilde{F}(x,c):x\in\Omega\cap\mathcal{C}\right\}.$
Besides, in view of (15), we derive that $1/g(x^{\star})\in\arg\max\{\widetilde{F}(x^{\star},c):c\in\mathbb{R}\}$, which along with (16) indicates that $(x^{\star},1/g(x^{\star}))$ is a global min-max point of (1.7), according to Definition 2.3.
Conversely, if $(x^{\star},c_{\star})$ is a global min-max point of (9), then we derive from Definition 2.3 that $x^{\star}$ satisfies (16) . Thus, we conclude that $x^{\star}\in\arg\min\{F(x):x\in\Omega\cap\mathcal{C}\}$ from (16) and (15). This completes the proof.
We next establish the connection between the local minimizes of problem (1) and the local min-max points of problem (9).
Proposition 3.3.
If $x^{\star}$ is a local minimizer of problem (1), then $(x^{\star},1/g(x^{\star}))$ is a local min-max point of problem (9). Conversely, if $(x^{\star},c_{\star})$ is a local min-max point of problem (9), then $x^{\star}$ is a local minimizer of problem (1).
Proof 3.4.
We first prove that the statement that $(x^{\star},1/g(x^{\star}))$ is a local min-max point of problem (9) follows from the $x^{\star}$ being a local minimizer of problem (1). Since (15) indicates that $1/g(x^{\star})$ is a global maximizer of $\widetilde{F}(x^{\star},\cdot)$, it suffices to show that, for all $\epsilon>0$, the $x^{\star}$ is always a local minimizer of
(17) $\varphi_{\epsilon}(x):=\max\left\{\widetilde{F}(x,c):c\in\mathbb{R},~{}|c-1/g(x^{\star})|\leq\epsilon\right\},$
relative to $\Omega\cap\mathcal{C}$ according to Definition 2.4. Let $\epsilon>0$. Firstly, due to the continuity of $1/g$ around $x^{\star}\in\Omega$, there exists $\delta_{\epsilon}>0$ such that $|1/g(x)-1/g(x^{\star})|\leq\epsilon$ for all $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})$. In view of this, we deduce from the definition (17) that
(18) $\widetilde{F}(x,1/g(x))\leq\varphi_{\varepsilon}(x)\text{,~{}~{} for all }x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})\text{. }$
Secondly, since $x^{\star}$ is a local minimizer of (1), there exists $\delta_{\epsilon}^{\prime}\in(0,\delta_{\epsilon})$, such that
(19) $F(x^{\star})\leq F(x)\text{,~{}~{} for all }x\in\mathcal{B}(x^{\star},\delta_{\epsilon}^{\prime})\cap(\Omega\cap\mathcal{C})\text{. }$
Thus, it holds for all $x\in\mathcal{B}(x^{\star},\delta_{\epsilon}^{\prime})\cap(\Omega\cap\mathcal{C})$ that
(20) $\varphi_{\epsilon}(x^{\star})\stackrel{{\scriptstyle{(I)}}}{{\leq}}\max_{c\in\mathbb{R}}\widetilde{F}(x^{\star},c)\stackrel{{\scriptstyle{(II)}}}{{=}}F(x^{\star})\stackrel{{\scriptstyle{(III)}}}{{\leq}}F(x)\stackrel{{\scriptstyle{(IV)}}}{{=}}\widetilde{F}(x,1/g(x))\stackrel{{\scriptstyle{(V)}}}{{\leq}}\varphi_{\varepsilon}(x),$
where ${(I)}\sim{(V)}$ follow from (17), (15), (19), (15) and (18), respectively. According to (20), we deduce that $x^{\star}$ is a local minimizer of $\varphi_{\epsilon}$ for an arbitrary $\epsilon>0$.
Conversely, we let $(x^{\star},c_{\star})$ be a local min-max point of (9). Then, according to Definition 2.4, we know that $c_{\star}$ is a local maximizer of $\widetilde{F}(x^{\star},\cdot)$. Hence, we have $0=\nabla(\widetilde{F}(x^{\star},\cdot))(c_{\star})$, which leads to $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. In addition, also according to Definition 2.4, we know that $x^{\star}$ is a local minimizer of $\varphi_{\epsilon}$ with some $\epsilon>0$ relative to $\Omega\cap\mathcal{C}$. Thus, there exists some $\delta_{\epsilon}>0$ such that
(21) $\varphi_{\epsilon}(x^{\star})\leq\varphi_{\epsilon}(x),\text{\quad for all $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})$.}$
For the left side of (21), we have
(22) $F(x^{\star})=\widetilde{F}(x^{\star},c_{\star})\leq\varphi_{\epsilon}(x^{\star}),$
where the first equation follows from (15) and the fact that $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. As to the right side of (21), it is clear that
(23) $\varphi_{\epsilon}(x)\leq\max\left\{\widetilde{F}(x,c):c\in\mathbb{R}\right\}=F(x),\quad\text{for all $x\in\Omega\cap\mathcal{C}$.}$
Together with (22) and (23), we deduce from (21) that $F(x^{\star})\leq F(x)$ for any $x\in\mathcal{B}(x^{\star},\delta_{\epsilon})\cap(\Omega\cap\mathcal{C})$, which yields that $x^{\star}$ is a local minimizer of (1).
Now we study the relationship between stationary points of problems (1) and (9). To this end, for $(x,c)\in(\Omega\cap\mathcal{C})\times\mathbb{R}$, we let ^$\widehat{\partial}_{x}\widetilde{F}(x,c)$ and $\nabla_{c}\widetilde{F}(x,c)$ denote the partial Fréchet subdifferential of $\widetilde{F}$ with respect to $x$ and the partial derivative of $\widetilde{F}$ with respect to $c$, respectively.
Proposition 3.5.
If $0\in\widehat{\partial}F(x^{\star})$, then for $c_{\star}=1/g(x^{\star})$ it holds that
[…]
Conversely, if $(x^{\star},c_{\star})$ satisfies (24), then $0\in\widehat{\partial}F(x^{\star})$.
Proof 3.6.
Let $(x^{\star},c_{\star})\in(\Omega\cap\mathcal{C})\times\mathbb{R}$ Due to the continuity of $g$ around $x^{\star}\in\Omega$, there exists $\delta>0$ such that $g(x)>0$ holds for all $x\in\mathcal{B}(x^{\star},\delta)$. Then, we have
(25) ^$\widehat{\partial}_{x}\widetilde{F}(x^{\star},c_{\star})=\widehat{\partial}\left(f/g+h+\iota_{\mathcal{C}}-\varphi\right)(x^{\star})=\widehat{\partial}(F-\varphi)(x^{\star}),$
where $\varphi(x):=\frac{f(x)}{g(x)}-2c_{\star}f(x)+c_{\star}^{2}f(x)g(x)$, and equally, we have
(26) $\varphi(x)=f(x)g(x)\left(\frac{1}{g(x)}-c_{\star}\right)^{2},\text{~{}~{}~{} for all $x\in\mathcal{B}(x^{\star},\delta)$}.$
We first show that (24) holds if $0\in\widehat{\partial}F(x^{\star})$ and $c_{\star}=1/g(x^{\star})$. Firstly, since $1/g(x^{\star})\in\arg\max\{\widetilde{F}(x^{\star},c):c\in\mathbb{R}\}$ (see (15)) and $\widetilde{F}(x^{\star},\cdot)$ is differentiable, we derive that $0=\nabla_{c}\widetilde{F}(x^{\star},c_{\star})$ for the $c_{\star}=1/g(x^{\star})$. Secondly, according to (26) and $c_{\star}=1/g(x^{\star})$, we have $|\varphi(x)-\varphi(x^{\star})|=f(x)(g(x))^{-1}(g(x^{\star}))^{-2}|g(x)-g(x^{\star})|^{2}$. This implies that $\varphi$ is differentiable at $x^{\star}$ with $\nabla\varphi\left(x^{\star}\right)=0$, since the real-valued convex function $g$ is Lipschitz continuous on $\mathcal{B}(x^{\star},\delta)$. Therefore, $0\in\widehat{\partial}F(x^{\star})$ yields $0\in\widehat{\partial}(F-\varphi)(x^{\star})$, which together with (25) implies that $0\in\widehat{\partial}_{x}\widetilde{F}(x^{\star},c_{\star})$.
Conversely, we suppose that (24) holds. Then $0=\nabla_{c}\widetilde{F}(x^{\star},c_{\star})$ implies that $f(x^{\star})c_{\star}=f(x^{\star})/g(x^{\star})$. Plugging this into (26) and noting that $f$ and $g$ are non-negative, we deduce that $\varphi(x^{\star})=0$ is the minimum of $\varphi$ on $x\in\mathcal{B}(x^{\star},\delta)$. Hence, by invoking the generalized Fermat’s Rule, we derive $0\in\widehat{\partial}\varphi(x^{\star})$. With the help of [30, Theorem 3.1], we have
(27) ^$\widehat{\partial}(F-\varphi)(x^{\star})\subseteq\bigcap_{v^{\star}\in\widehat{\partial}\varphi(x^{\star})}(\widehat{\partial}F(x^{\star})-v^{\star})\subseteq\widehat{\partial}F(x^{\star})-\{0\}.$
Together with the first relation of (24), the equation (25), and the relation (27), we finally conclude that $0\in\widehat{\partial}F(x^{\star})$. This completes the proof.
With the help of Proposition 3.5, we present a useful first-order necessary condition for problem (1).
Corollary 3.7.
If $0\in\widehat{\partial}F(x^{\star})$, then for $c_{\star}=1/g(x^{\star})$ it holds that
(28) $0\in\widehat{\partial}(c_{\star}f+\iota_{\mathcal{C}})\left(x^{\star}\right)-c_{\star}^{2}f(x^{\star})\partial g\left(x^{\star}\right)+\nabla h_{1}\left(x^{\star}\right)-\partial h_{2}\left(x^{\star}\right).$
Moreover, if $g$ and $h_{2}$ are differentiable at $x^{\star}$, then $0\in\widehat{\partial}F(x^{\star})$ is equivalent to (28) with $c_{\star}=1/g(x^{\star})$.
Proof 3.8.
Note that $\varphi_{2}:=2c_{\star}f-c_{\star}^{2}fg-c_{\star}f+c_{\star}^{2}f(x^{\star})g$ is differentiable at $x^{\star}$. Precisely, we have $\nabla\varphi_{2}(x^{\star})=0$, since
$\lim_{x\to x^{\star}}\frac{|\varphi_{2}(x)-\varphi_{2}(x^{\star})|}{\|x-x^{\star}\|_{2}}=\lim_{x\to x^{\star}}\frac{c_{\star}^{2}|f(x)-f(x^{\star})||g(x^{\star})-g(x)|}{\|x-x^{\star}\|_{2}}=0,$
where the last equality holds, owing to $x^{\star}\in\mathrm{dom}(\widehat{\partial}F)\subseteq\Omega$, and the local Lipschitz continuity of $f$ and $g$ on $\Omega$. Hence, for $c_{\star}=1/g(x^{\star})$, it holds that
(29) ^$\widehat{\partial}_{x}\widetilde{F}(x^{\star},c_{\star})\overset{{(I)}}{=}\widehat{\partial}(2c_{\star}f-c_{\star}^{2}fg+h+\iota_{\mathcal{C}})(x^{\star})$
=$\overset{{(II)}}{=}\widehat{\partial}(c_{\star}f-c_{\star}^{2}f(x^{\star})g-h_{2}+\iota_{\mathcal{C}})(x^{\star})+\nabla\varphi_{2}(x^{\star})+\nabla h_{1}(x^{\star})$
$\overset{{(III)}}{\subseteq}\widehat{\partial}(c_{\star}f+\iota_{\mathcal{C}})\left(x^{\star}\right)-c_{\star}^{2}f(x^{\star})\partial g\left(x^{\star}\right)+\nabla h_{1}\left(x^{\star}\right)-\partial h_{2}\left(x^{\star}\right),$
where ${(I)}$ follows from the definition of $\widetilde{F}$ (see (13)) and the open $\Omega$, ${(II)}$ holds because $\varphi_{2}$ and $h_{1}$ are differentiable at $x^{\star}$, and ${(III)}$ holds thanks to [30, Theorem 3.1], $c_{\star}^{2}f(x^{\star})\geq 0$, and the real-valued convexity of the functions $g$ and $h_{2}$. Especially, ${(III)}$ in (29) becomes an equality when $g$ and $h_{2}$ are differentiable at $x^{\star}$. Invoking (29) and Proposition 3.5, we then obtain the desired results.
Inspired by Corollary 3.7, we introduce the definition of a critical point of $F$.
Definition 3.9.
We say that $x^{\star}\in\Omega$ is a critical point of $F$, if (28) holds with $c_{\star}=1/g(x^{\star})$.
To close this section, we make some remarks regarding Definition 3.9. In view of Corollary 3.7 and the generalized Fermat’s Rule, a local minimizer of $F$ must be a critical point of $F$. In particular, when both $g$ and $h_{2}$ are differentiable at $x^{\star}$, according to Corollary 3.7 we see that the statement that $x^{\star}$ is a critical point of $F$ coincides with $0\in\widehat{\partial}F(x^{\star})$.
4 An alternating maximization proximal descent algorithm
In this section, we propose an alternating maximization proximal descent algorithm (AMPDA) for solving the min-max reformulation (9) and establish its subsequential convergence to a critical point of problem (1) under a mild assumption.
4.1 AMPDA and its interpretations
In this subsection, we present the AMPDA and then make some interpretations on this algorithm. To this end, we first introduce the notion of proximal operators. For a proper closed function $\varphi:\mathbb{R}^{n}\to(-\infty,+\infty]$, the proximal operator of $\varphi$ at $x\in\mathbb{R}^{n}$ is denoted by $\mathrm{prox}_{\varphi}(x)$ and defined as
(30) $\mathrm{prox}_{\varphi}(x):=\arg\min\left\{\varphi(z)+\frac{1}{2}\|x-z\|^{2}_{2}:z\in\mathbb{R}^{n}\right\}.$
The proximal operator $\mathrm{prox}_{\varphi}$ takes an $x\in\mathbb{R}^{n}$ and maps it into a subset of $\mathbb{R}^{n}$, which may be empty, a singleton or a subset with multiple vectors. If for a given $x\in\mathbb{R}^{n}$, the function $z\to\varphi(z)+\frac{1}{2}\|x-z\|^{2}_{2}$ is coercive, then $\mathrm{prox}_{\varphi}(x)$ is nonempty (see, for example, [4, Theorem 6.4]). Moreover, if $\varphi$ is convex, then $\mathrm{prox}_{\varphi}(x)$ is a singleton for all $x\in\mathbb{R}^{n}$. In particular, let $f_{\mathcal{C}}:=f+\iota_{\mathcal{C}}$, then the proximal operator $\mathrm{prox}_{f_{\mathcal{C}}}(x)$ is nonempty for all $x\in\mathbb{R}^{n}$ due to $f_{\mathcal{C}}(x)\geq 0$. Let $Q:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\to(-\infty,+\infty]$ be defined by
(31) $Q(x,y,z,c)=\iota_{\mathcal{C}}(x)+2cf(x)+c^{2}f(x)(g^{*}(y)-\left\langle{x},{y}\right\rangle)+h_{1}(x)+h_{2}^{*}(z)-\left\langle{x},{z}\right\rangle$
if $(y,z)\in\mathrm{dom}(g^{*})\times\mathrm{dom}(h_{2}^{*})$, and $Q(x,y,z,c)=+\infty$ otherwise. Let $F$ and $\widetilde{F}$ be defined by (12) and (13), respectively. Then, for all $x\in\Omega\cap\mathcal{C}$, $(y,z)\in\mathrm{dom}(g^{*})\times\mathrm{dom}(h_{2}^{*})$ and $c_{x}=1/g(x)$, it holds that
(32) $F(x)\overset{{(I)}}{=}\widetilde{F}(x,c_{x})\overset{{(II)}}{\leq}Q(x,y,z,c_{x}),$
where ${(I)}$ follows from (15), ${(II)}$ follows from the Fenchel-Young Inequality and the non-negativity of $c_{x}^{2}f(x)$, and ${(II)}$ becomes an equality when $(y,z)\in\partial g(x)\times\partial h_{2}(x)$. Now we present our AMPDA stated in Algorithm 1. Before proceeding, we provide some interpretations and observations on the AMPDA. This algorithm can be interpreted as an alternating iteration method for solving the min-max reformulation (9) as illustrated below.
Algorithm 1 AMPDA for solving problem (1).
Step 0. Input $x^{0}\in\mathrm{dom}(F)$, $0<\underline{\alpha}\leq\overline{\alpha}$, $\sigma>0$, $10<\gamma<$ 1, and set $t\leftarrow 0$.
Step 1. Compute $c_{k}=1/g(x^{k})$.
Choose $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$.
Set $\alpha:=\widetilde{\alpha}_{k}\in[\underline{\alpha},\overline{\alpha}]$.
Step 2. Compute
(33) $\widehat{x}^{k}\in\mathrm{prox}_{\alpha c_{k}f_{\mathcal{C}}}\left(x^{k}-\alpha(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})\right).$
If $\widehat{x}^{k}\in\Omega$ and satisfies
(34) $Q(\widehat{x}^{k},y^{k},z^{k},1/g(\widehat{x}^{k}))+\frac{\sigma}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq F(x^{k});$
Then, set $x^{k+1}=\widehat{x}^{k}$ and go to Step 3;
Else, set $\alpha:=\alpha\gamma$ and repeat Step 2;
Step 3. Record $\alpha_{k}:=\alpha$, set $k\leftarrow k+1$, and go to Step 1.
First, for fixed $x\in\Omega\cap\mathcal{C}$, $\widetilde{F}(x,\cdot)$ is either a concave one-dimensional quadratic function or a constant function, which always admits a global maximizer at $c=1/g(x)$. Hence, given $x^{k}\in\Omega\cap\mathcal{C}$, AMPDA sets $c_{k}=1/g(x^{k})$ to maximize $\widetilde{F}(x^{k},\cdot)$ in Step 1.
Second, in Step 2 of the $k$-th iteration, AMPDA solves a proximal subproblem with respect to $f+\iota_{\mathcal{C}}$, which also requires to evaluate the gradient of $h_{1}$, the subgradients of $g$ and $h_{2}$ at $x^{k}$. As we will see below, this can be viewed as applying one step of the majorization-minimization scheme to the minimization problem $\min\{\widetilde{F}(x,c_{k}):x\in\mathbb{R}^{n}\}$. Since $f$ and $\nabla h_{1}$ are locally Lipschitz continuous on $\Omega$, there exists $\delta_{k}>0$ such that $f$ and $\nabla h_{1}$ are Lipschitz continuous on the neighborhood $\mathcal{B}(x^{k},\delta_{k})\subseteq\Omega$. Let $L_{f,k}>0$ and $L_{\nabla h_{1},k}>0$ be the associated Lipschitz moduli of $f$ and $\nabla h_{1}$ respectively. In the majorization stage, we introduce the surrogate function $\mathcal{M}(\cdot\mid x^{k},c_{k})$ of $\widetilde{F}(\cdot,c_{k})$ over $\mathcal{B}(x^{k},\delta_{k})$, which is in the form of
(35) $\mathcal{M}(x\mid x^{k},c_{k})=(2c_{k}-c_{k}^{2}g(x^{k}))f(x)+\left\langle\nabla h_{1}(x^{k})-c_{k}^{2}f(x^{k})y^{k}-z^{k},x-x^{k}\right\rangle\\ +\left(c_{k}^{2}L_{f,k}\|y^{k}\|_{2}+\frac{L_{\nabla h_{1},k}}{2}\right)\|x-x^{k}\|_{2}^{2}+h(x^{k})+\iota_{\mathcal{C}}(x).$
Clearly, $\mathcal{M}(x^{k}\mid x^{k},c_{k})=\widetilde{F}(x^{k},c_{k})$. Moreover, for all $x\in\mathcal{B}(x^{k},\delta_{k})$, we have
$\widetilde{F}(x,c_{k})\overset{{(I)}}{=}\iota_{\mathcal{C}}(x)+f(x)\left(2c_{k}-c_{k}^{2}g(x)\right)+h(x)$
$\overset{{(II)}}{\leq}\iota_{\mathcal{C}}(x)+f(x)\left(2c_{k}-c_{k}^{2}\left(g(x^{k})+\left\langle y^{k},x-x^{k}\right\rangle\right)\right)+h(x)$
$=\iota_{\mathcal{C}}(x)+\left(2c_{k}-c_{k}^{2}g(x^{k})\right)f(x)-\left\langle c_{k}^{2}f(x^{k})y^{k},x-x^{k}\right\rangle+h(x)$
$\quad+\left\langle c_{k}^{2}y^{k},x-x^{k}\right\rangle(f(x^{k})-f(x))$
$\overset{{(III)}}{\leq}\iota_{\mathcal{C}}(x)+\left(2c_{k}-c_{k}^{2}g(x^{k})\right)f(x)-\left\langle c_{k}^{2}f(x^{k})y^{k},x-x^{k}\right\rangle+h(x)$
$\qquad+c_{k}^{2}L_{f,k}\|y^{k}\|_{2}\|x-x^{k}\|_{2}^{2}\overset{{(IV)}}{\leq}\mathcal{M}(x\mid x^{k},c_{k}),$
where ${(I)}$ follows from the definition of $\widetilde{F}$ (see (13)), ${(II)}$ follows from $f\geq 0$ and $y^{k}\in\partial g(x^{k})$, ${(III)}$ follows from the Lipschitz continuity of $f$ and the Cauchy-Schwarz Inequality, and ${(IV)}$ holds since the Lipschitz continuity of $\nabla h_{1}$ on $\mathcal{B}(x^{k},\delta_{x})$ and $z^{k}\in\partial h_{2}(x^{k})$ imply that
$h(x)=h_{1}(x)-h_{2}(x)$
$\leq h_{1}(x^{k})-h_{2}(x^{k})+\left\langle\nabla h_{1}(x^{k})-z^{k},x-x^{k}\right\rangle+\frac{L_{\nabla h_{1},k}}{2}\|x-x^{k}\|_{2}^{2}$
$=h(x^{k})+\left\langle\nabla h_{1}(x^{k})-z^{k},x-x^{k}\right\rangle+\frac{L_{\nabla h_{1},k}}{2}\|x-x^{k}\|_{2}^{2}.$
Next, in the minimization stage one naturally updates $x$ by solving $\min\{\mathcal{M}(x\mid x^{k},c_{k}):x\in\mathbb{R}^{n}\}$. However, as the Lipschitz moduli $L_{f,k}$ and $L_{\nabla h_{1},k}$ are generally difficult to estimate, we use $1/(2\alpha)$ with $\alpha>0$ to replace the unknown term $c_{k}^{2}L_{f,k}\|y^{k}\|_{2}+\frac{L_{\nabla h_{1},k}}{2}$ in $\mathcal{M}(x\mid x^{k},c_{k})$, where $\alpha$ is determined by a backtracking line search procedure. Invoking this, (35) and the fact that $c_{k}=1/g(x^{k})$, we actually in the minimization stage construct $x^{k+1}$ via
(36) $x^{k+1}\in\arg\min\Big{\{}c_{k}f(x)+\left\langle\nabla h_{1}(x^{k})-c_{k}^{2}f(x^{k})y^{k}-z^{k},x-x^{k}\right\rangle\\ +\frac{1}{2\alpha}\|x-x^{k}\|_{2}^{2}:x\in\mathcal{C}\Big{\}}.$
It is obvious that (36) is exactly equivalent to (33) in Step 2 of AMPDA.
Third, the line search procedure is stopped once $\widehat{x}^{k}\in\Omega$ and (34) is satisfied. Note that $(y^{k},z^{k})\in\mathrm{dom}(g^{*})\times\mathrm{dom}(h_{2}^{*})$ holds for all $k\in\mathbb{N}$, owing to $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$. Invoking this and (32), we immediately see that the condition (34) is generally stronger than the sufficient descent condition of $F$ at $\widehat{x}^{k}$, that is,
(37) $F(\widehat{x}^{k})+\frac{\sigma}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq F(x^{k}).$
Although the subsequential convergence can be still guaranteed if the terminating condition (34) of the line search scheme in AMPDA is replaced by (37), it is generally difficult to show the convergence of the full solution sequence generated by this altered algorithm. In contrast, we can not only establish the subsequential convergence but also the convergence of the entire solution sequence generated by the AMPDA equipped with (34) under the KL assumption.
4.2 Convergence analysis for the AMPDA
In this subsection, we conduct convergence analysis for the AMPDA. To this end, we first introduce the following basic assumption concerning its initial point.
Assumption 2
The level set $\mathcal{X}:=\{x\in\mathrm{dom}(F):F(x)\leq F(x^{0})\}$ is compact.
Note that the boundedness of the level set associated with the extended objective is a very standard assumption in nonconvex optimization, while its closeness automatically holds if the extended objective is closed. Nevertheless, the $F$ defined in (12) may be not closed, which necessitates the closedness condition in 2. In view of [44, Proposition 4.4] and the closedness of $h$ and $\mathcal{C}$, we see that $F$ is closed on $\mathbb{R}^{n}$ if $f$ and $g$ do not attain zero simultaneously. In the case where $f$ and $g$ can attain zero simultaneously, we can still guarantee the closedness of $\mathcal{X}$ via appropriately selecting $x^{0}$, as shown in the following proposition.
Proposition 4.1.
Suppose that $\mathcal{O}:=\{x\in\mathbb{R}^{n}:f(x)=g(x)=0\}\neq\emptyset$. Then, $\mathcal{X}$ is closed, if $x^{0}\in\mathrm{dom}(F)$ satisfies
(38) $F(x^{0})<\inf\left\{\mathop{\lim\inf}_{z\to x}F(z):x\in\mathcal{O}\right\}.$
Proof 4.2.
We shall show that $\mathcal{X}$ is closed by verifying that each accumulation point of $\mathcal{X}$ belongs to $\mathcal{X}$ when (38) is satisfied. Let $x^{\star}$ be an accumulation point of $\mathcal{X}$ and the sequence $\{x^{k}:k\in\mathbb{N}\}\subseteq\mathcal{X}$ converges to $x^{\star}$. For any $k\in\mathbb{N}$, we derive from $x^{k}\in\mathcal{X}$ that $h(x^{k})+f(x^{k})/g(x^{k})\leq F(x^{0})$. This together with $g(x^{k})>0$ yields that, for any $k\in\mathbb{N}$,
(39) $f(x^{\star})+h(x^{k})g(x^{\star})$ $\leq F(x^{0})g(x^{k})+h(x^{k})(g(x^{\star})-g(x^{k}))$
$\leq F(x^{0})g(x^{k})+\sup\{|h(x^{k})|:k\in\mathbb{N}\}|g(x^{\star})-g(x^{k})|.$
Firstly, we have $\inf\{h(x^{k}):k\in\mathbb{N}\}>-\infty$ from $x^{k}\to x^{\star}$ and the closedness of $h$. Secondly, $\sup\{h(x^{k}):k\in\mathbb{N}\}=\sup\{F(x^{k})-f(x^{k})/g(x^{k}):k\in\mathbb{N}\}\leq F(x^{0})$ follows from $x^{k}\in\mathcal{X}$ for each $k\in\mathbb{N}$, and the fact that $F\leq F(x^{0})$ and $f/g\geq 0$ holds on $\mathcal{X}$. In view of these two reasons, we deduce that $\sup\{|h(x^{k})|:k\in\mathbb{N}\}<+\infty$. Therefore, by passing to the limit on the both sides of (39) with $k\to\infty$, we derive that
(40) $f(x^{\star})+h(x^{\star})g(x^{\star})\leq F(x^{0})g(x^{\star}),$
owing to the closed $f$, the closed $g(x^{\star})h$, and the continuous $g$. From (40) and the closedness of $\mathcal{C}$, one can derive $x^{\star}\in\mathcal{X}$ if $g(x^{\star})>0$. In fact, $g(x^{\star})>0$ is guaranteed when $x^{0}\in\mathrm{dom}(F)$ satisfies (38). Assume on the contrary that (38) holds but $g(x^{\star})=0$, which forces $f(x^{\star})=0$ according to (40) and $f\geq 0$. In other words, $g(x^{\star})=0$ leads to $x^{\star}\in\mathcal{O}$. This, along with (38), implies that $F(x^{0})<\mathop{\lim\inf}_{k\to\infty}F(x^{k})$, which contradicts the fact that $\{x^{k}:k\in\mathbb{N}\}\subseteq\mathcal{X}$.
We remark that for both problems (2) and (3), the zero vector is the unique point at which both $f$ and $g$ vanish. In view of Proposition 4.1 and these two problems, we deduce that if $x^{0}\in\mathbb{R}^{n}$ satisfies $F(x^{0})<\mathop{\lim\inf}_{z\to 0}F(z)=1+\frac{\lambda}{2}\|b-\mathcal{T}_{\mu}(b)\|_{2}^{2}$, then the level set $\mathcal{X}$ is closed. This together with the boundedness of the constraint $\{x\in\mathbb{R}^{n}:\underline{x}\leq x\leq\overline{x}\}$ shows that 2 is fulfilled for problems (2) and (3) if the initial point $x^{0}$ is properly chosen such that
(41) $F(x^{0})<1+\frac{\lambda}{2}\|b-\mathcal{T}_{\mu}(b)\|_{2}^{2}.$
Now with the help of 2, we can establish the following technical lemma, which will be frequently used in our subsequent convergence analysis.
Lemma 4.3.
Suppose that 2 holds. Then, we have $m_{g}:=\inf\{g(x):x\in\mathcal{X}\}>0$. Moreover, there exists $\Delta>0$ such that the following statements hold with the set $\mathcal{X}_{\Delta}$ defined by
$\mathcal{X}_{\Delta}:=\{x\in\mathbb{R}^{n}:\text{$\|x-\bar{x}\|_{2}\leq\Delta$for some $\bar{x}\in\mathcal{X}$}\}.$
  1. (i)
    For any $x\in\mathcal{X}_{\Delta}$, there holds $g(x)\geq m_{g}/2$.
  2. (ii)
    $f$, $g$, $1/g$ and $\nabla h_{1}$ are globally Lipschitz continuous on $\mathcal{X}_{\Delta}$.
  3. (iii)
    The quantities defined below are finite.
    $M_{\partial g}:=\sup\{\|y\|_{2}:y\in\partial g(x)\text{~{}for some $x\in\mathcal{X}$}\},$
    $M_{f/g^{2}}:=\sup\{f(x)/g^{2}(x):x\in\mathcal{X}_{\Delta}\},~{}~{}~{}~{}~{}M_{g}:=\sup\{g(x):x\in\mathcal{X}\}.$
Proof 4.4.
Since the continuous $g$ is strictly positive on the compact $\mathcal{X}$, we deduce that $m_{g}>0$ and the statement (i) holds. Then, $f$, $g$, and $\nabla h_{1}$ are globally Lipschitz continuous on $\mathcal{X}_{\Delta}$, since they are locally Lipschitz continuous around each point in the compact $\mathcal{X}_{\Delta}\subseteq\Omega$ ([13, Chapter 1, 7.5 Exercise (c)]). This, along with the statement (i), implies that the statement (ii) holds. Lastly, $M_{\partial g}$, $M_{f/g^{2}}$, and $M_{g}$ are finite due to the boundedness of $\cup_{x\in\mathcal{X}}\partial g(x)$, continuity of $f/g^{2}$ on the compact $\mathcal{X}_{\Delta}$, and the continuity of $g$ on the compact $\mathcal{X}$, respectively. This completes the proof.
Next, we prove the well-definedness of AMPDA, which also implies its sufficient descent property.
Proposition 4.5.
Suppose that 2 holds. Then the AMPDA is well-defined. Specifically, there exists $\underline{\alpha}_{\sigma}>0$ such that, for all $k\in\mathbb{N}$, Step 2 of AMPDA terminates at some $\alpha_{k}\geq\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. Moreover, the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA falls into the level set $\mathcal{X}\subseteq\Omega\cap\mathcal{C}$ and satisfies
(42) $F(x^{k+1})+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|^{2}_{2}\leq F(x^{k}).$
Proof 4.6.
We prove this proposition by induction. It is clear that $x^{0}\in\mathcal{X}$ and we assume that $x^{k}\in\mathcal{X}$ for some $k\geq 0$.
Let $\mathcal{X}_{\Delta}$ be given as Lemma 4.3. We first show that there exists $\underline{\alpha}_{\Delta}>0$ such that $\widehat{x}^{k}$ falls into the set $\mathcal{X}_{\Delta}\cap\mathcal{C}$, whenever $\alpha\in(0,\underline{\alpha}_{\Delta}]$. From (33), we know that
(43) $\alpha c_{k}f(\widehat{x}^{k})+\frac{1}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}+\left\langle\widehat{x}^{k}-x^{k},\alpha(\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k})\right\rangle\\ \leq\alpha c_{k}f(x^{k}).$
The relation (43) implies that $\alpha c_{k}f(\widehat{x}^{k})+\frac{1}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}-\alpha\|\widehat{x}^{k}-x^{k}\|_{2}\|\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k}\|_{2}\leq\alpha c_{k}f(x^{k})$, which is a quadratic inequality of the term $\|\widehat{x}^{k}-x^{k}\|_{2}$. In view of this, $c_{k}=1/g(x^{k})>0$ and $f(\widehat{x}^{k})\geq 0$, we deduce that
(44) $\|\widehat{x}^{k}-x^{k}\|_{2}\leq\alpha|g(x^{k})|\left\|\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k}\right\|_{2}\\ +\sqrt{\alpha^{2}|g(x^{k})|^{2}\|\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k}\|_{2}^{2}+2\alpha f(x^{k})}.$
Since the compact $\mathcal{X}\subseteq\Omega$ implies that $\sup\{|g(x)|\|\nabla h_{1}(x)-z-c^{2}f(x)y\|_{2}:x\in\mathcal{X},(y,z)\in\partial g(x)\times\partial h_{2}(x),c=1/g(x)\}<+\infty$ and $\sup\{f(x):x\in\mathcal{X}\}<+\infty$, the relation (44) along with $x^{k}\in\mathcal{X}$ implies that the term $\|\widehat{x}^{k}-x^{k}\|_{2}$ can be narrow down by the parameter $\alpha$. Therefore, there exists some $\underline{\alpha}_{\Delta}>0$ such that
(45) $\|\widehat{x}^{k}-x^{k}\|_{2}\leq\Delta/2,\text{~{}~{}~{}~{}~{}for any $\alpha\in(0,\underline{\alpha}_{\Delta}]$}.$
This, together with $x^{k}\in\mathcal{X}$ and (33), yields that $\widehat{x}^{k}\in\mathcal{X}_{\Delta}\cap\mathcal{C}$ when $\alpha\in(0,\underline{\alpha}_{\Delta}]$.
Next, we shall show that there exists $\underline{\alpha}_{\sigma}\in(0,\underline{\alpha}_{\Delta}]$ such that $\widehat{x}^{k}\in\Omega$ and satisfies (34) stated in Step 2 of AMPDA whenever $\alpha\in(0,\underline{\alpha}_{\sigma}]$. Let $\alpha\in(0,\underline{\alpha}_{\Delta}]$. Clearly, with the help of Lemma 4.3 (i), the statement that $\widehat{x}^{k}\in\Omega$ holds due to $\widehat{x}^{k}\in\mathcal{X}_{\Delta}$. Besides, from (45), we have $\widehat{x}^{k}\in\mathcal{B}(x^{k},\Delta)$. This, together with the Lipschitz continuity of $\nabla h_{1}$ on $\mathcal{B}(x^{k},\Delta)\subseteq\mathcal{X}_{\Delta}$ (see Lemma 4.3 (ii)), implies that
(46) $h_{1}(\widehat{x}^{k})\leq h_{1}(x^{k})+\left\langle{\nabla h_{1}(x^{k})},{\widehat{x}^{k}-x^{k}}\right\rangle+\frac{L_{\nabla h_{1}}}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2},$
where $L_{\nabla h_{1}}>0$ denotes the Lipschitz modulus of $\nabla h_{1}$ on $\mathcal{X}_{\Delta}$. Multiplying $1/\alpha$ on the both sides of (43) and summing (46), we derive that
(47) $c_{k}f(\widehat{x}^{k})+\left\langle x^{k}-\widehat{x}^{k},z^{k}+c^{2}_{k}f(x^{k})y^{k}\right\rangle+h_{1}(\widehat{x}^{k})\\ +c_{k}\left(\frac{1}{2\alpha}-\frac{L_{\nabla h_{1}}}{2}\right)\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq c_{k}f(x^{k})+h_{1}(x^{k}).$
In addition, by letting $\hat{c}=1/g(\widehat{x}^{k})$, the direct computation yields that
(48) $c_{k}f(\widehat{x}^{k})+\left\langle x^{k}-\widehat{x}^{k},c^{2}_{k}f(x^{k})y^{k}\right\rangle$
$\qquad-\Big{(}\hat{c}^{2}f(\widehat{x}^{k})\left(2g(\widehat{x}^{k})-g(x^{k})\right)+\left\langle x^{k}-\widehat{x}^{k},\hat{c}^{2}f(\widehat{x}^{k})y^{k}\right\rangle\Big{)}$
$=c_{k}\hat{c}^{2}f(\widehat{x}^{k})\left(g(\widehat{x}^{k})-g(x^{k})\right)^{2}+\left\langle x^{k}-\widehat{x}^{k},\left(c^{2}_{k}f(x^{k})-\hat{c}^{2}f(\widehat{x}^{k})\right)y^{k}\right\rangle$
$\stackrel{{\scriptstyle{(I)}}}{{\geq}}-\left(c_{k}\hat{c}^{2}f(\widehat{x}^{k})L_{g}^{2}+L_{f/g^{2}}\|y^{k}\|_{2}\right)\|x^{k}-\widehat{x}^{k}\|_{2}^{2}$
$\stackrel{{\scriptstyle{(II)}}}{{\geq}}-c_{k}\left(M_{f/g^{2}}L_{g}^{2}+L_{f/g^{2}}M_{\partial g}M_{g}\right)\|x^{k}-\widehat{x}^{k}\|_{2}^{2},$
where ${(I)}$ follows from the Cauchy-Schwarz Inequality and Lemma 4.3 (ii) with the positive $L_{g}$ and $L_{f/g^{2}}$ being the Lipschitz moduli of $g$ and $f/g^{2}$ on $\mathcal{X}_{\Delta}$, respectively, and ${(II)}$ follows from Lemma 4.3 (iii). Combining (47) and (48), we derive that
(49) $\hat{c}^{2}f(\widehat{x}^{k})\left(2g\left(\widehat{x}^{k}\right)-g(x^{k})\right)+\left\langle x^{k}-\widehat{x}^{k},z^{k}+\hat{c}^{2}f(\widehat{x}^{k})y^{k}\right\rangle+h_{1}(\widehat{x}^{k})\\ +c_{k}\left(\frac{1}{2\alpha}-\frac{L_{\nabla h_{1}}}{2}-M_{f/g^{2}}L_{g}^{2}-L_{f/g^{2}}M_{\partial g}M_{g}\right)\left\|\widehat{x}^{k}-x^{k}\right\|_{2}^{2}\leq c_{k}f(x^{k})+h_{1}(x^{k}).$
Recalling $\widehat{x}^{k}\in\mathcal{X}_{\Delta}\cap\mathcal{C}$, and invoking (31), $\left\langle{x^{k}},{y^{k}}\right\rangle=g(x^{k})+g^{*}(y^{k})$ and $\left\langle{x^{k}},{z^{k}}\right\rangle=h_{2}(z^{k})+h_{2}^{*}(z^{k})$, one can verify that (49) is equivalent to
$Q(\widehat{x}^{k},y^{k},z^{k},1/g(\widehat{x}^{k}))\\ +c_{k}\left(\frac{1}{2\alpha}-\frac{L_{\nabla h_{1}}}{2}-M_{f/g^{2}}L_{g}^{2}-L_{f/g^{2}}M_{\partial g}M_{g}\right)\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq F(x^{k}).$
Thus, the $\widehat{x}^{k}$ given by (33) falls into $\mathcal{X}_{\Delta}\cap\mathcal{C}\subseteq\Omega\cap\mathcal{C}$, and satisfies (34), if $\alpha\in(0,\underline{\alpha}_{\sigma}]$, where $\underline{\alpha}_{\sigma}>0$ is given by
$\underline{\alpha}_{\sigma}:=\min\left\{\underline{\alpha}_{\Delta},\left(L_{\nabla h_{1}}+2M_{f/g^{2}}L_{g}^{2}+\left(2L_{f/g^{2}}M_{\partial g}+\sigma\right)M_{g}\right)^{-1}\right\}.$
This implies that Step 2 of AMPDA terminates at some $\alpha_{k}\geq\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$ in the $k$-th iteration.
Furthermore, according to Step 2 of AMPDA in the $k$-th iteration, the derived $x^{k+1}\in\Omega\cap\mathcal{C}$ satisfies
(50) $Q(x^{k+1},y^{k},z^{k},1/g(x^{k+1}))+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|_{2}^{2}\leq F(x^{k}),$
which, together with (32), indicates (42). In view of (42) and $x^{k}\in\mathcal{X}$, we obtain $x^{k+1}\in\mathcal{X}$. Consequently, we conclude this proposition by induction.
Proposition 4.5 and the inequality (42) immediately lead to the next corollary.
Corollary 4.7.
Suppose that 2 holds. Then, the $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ generated by AMPDA is bounded. Besides, the limit $F_{\infty}:=\lim_{k\to\infty}F(x^{k})$ exists, and it holds that
(51) $\lim_{k\to\infty}\|x^{k+1}-x^{k}\|_{2}=0.$
Proof 4.8.
Firstly, since $x^{k}\in\mathcal{X}$ and the functions $g$ and $h_{2}$ are real-valued convex, we derive $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ is bounded due to $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$ for all $k\in\mathbb{N}$, $c_{k}=1/g(x^{k})$ and the compact $\mathcal{X}\subseteq\Omega$. Secondly, since $F$ is continuous on $\Omega\cap\mathcal{C}\supseteq\mathcal{X}$, the function $F$ is bounded below on the compact $\mathcal{X}$, which together with (42) indicates that the limit $F_{\infty}:=\lim_{k\to\infty}F(x^{k})$ exists. Thirdly, the equation (51) is derived by passing to the limit on the both sides of (42) with $k\to\infty$.
Now we are ready to prove the subsequential convergence of the proposed AMPDA.
Theorem 4.9.
Suppose that 2 holds. Then any accumulation point of the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA falls into the level set $\mathcal{X}$, and is a critical point of $F$.
Proof 4.10.
Let $x^{\star}$ be an accumulation point of $\{x^{k}:k\in\mathbb{N}\}$. In view of $\{x^{k}:k\in\mathbb{N}\}\subseteq\mathcal{X}$ (see Proposition 4.5), we derive from the compactness of $\mathcal{X}$ that $x^{\star}\in\mathcal{X}$. Next, we shall show that $x^{\star}$ is a critical point of $F$. Let $\mathcal{K}\subseteq\mathbb{N}$ be a subsequence, such that $\{x^{k}:k\in\mathcal{K}\}$ converges to the $x^{\star}$. Invoking Proposition 4.5 and (33), we have $x^{k+1}\in\mathrm{prox}_{\alpha_{k}c_{k}f_{\mathcal{C}}}\left(x^{k}-\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})\right)$ for some $\alpha_{k}\geq\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. Hence, for any $(x,k)\in\mathcal{C}\times\mathcal{K}$, it holds that
$\alpha_{k}c_{k}f(x^{k+1})+\frac{1}{2}\|x^{k+1}-x^{k}\|_{2}^{2}+\left\langle{x^{k+1}-x^{k}},{\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})}\right\rangle\\ \leq\alpha_{k}c_{k}f(x)+\frac{1}{2}\|x-x^{k}\|_{2}^{2}+\left\langle{x-x^{k}},{\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})}\right\rangle.$
This, together with the fact that $\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma\leq\alpha_{k}\leq\overline{\alpha}$, leads to
(52) $c_{k}f(x^{k+1})+\frac{1}{2\overline{\alpha}}\|x^{k+1}-x^{k}\|_{2}^{2}+\left\langle{x^{k+1}-x^{k}},{\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k}}\right\rangle\\ \leq c_{k}f(x)+\frac{1}{2\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma}\|x-x^{k}\|_{2}^{2}+\left\langle{x-x^{k}},{\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k}}\right\rangle.$
Since $\{(y^{k},z^{k}):k\in\mathcal{K}\}$ is bounded (see Corollary 4.7) and $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$ for all $k\in\mathbb{N}$, we deduce that some subsequence $\mathcal{K}_{2}\subseteq\mathcal{K}$ such that $\{(y^{k},z^{k}):k\in\mathcal{K}_{2}\}$ converges to some $(y^{\star},z^{\star})\in\partial g(x^{\star})\times\partial h_{2}(x^{\star})$. In view of (51) and the continuity of $f$ and $1/g$ on $\mathcal{X}$, we pass to the limit with $k\in\mathcal{K}_{2}$ and $k\to\infty$ on the both sides of (52), and then obtain that
(53) $c_{\star}f(x^{\star})\leq c_{\star}f(x)+\frac{\|x-x^{\star}\|_{2}^{2}}{2\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma}+\left\langle{x-x^{\star}},{\nabla h_{1}(x^{\star})-z^{\star}-c_{\star}^{2}f(x^{\star})y^{\star}}\right\rangle,$
for all $x\in\mathcal{C}$, where $c_{\star}=1/g(x^{\star})$. The relation (53) implies that
(54) $x^{\star}\in\arg\min\left\{c_{\star}f_{\mathcal{C}}(x)+\frac{1}{2\alpha}\|x-x^{\star}+\alpha(\nabla h_{1}(x^{\star})-z^{\star}-c_{\star}^{2}f(x^{\star})y^{\star})\|^{2}_{2}:x\in\mathbb{R}^{n}\right\}$
with $\alpha=\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. By invoking the generalized Fermat’s Rule, the relation (54) yields (28), which together with the fact $x^{\star}\in\mathcal{X}\subseteq\mathrm{dom}(F)$ indicates that $x^{\star}$ is a critical point of $F$ according to Definition 3.9. This completes the proof.
To close this subsection, we show that the proximal step (33) in AMPDA relates to the update formula (6) that is used in existing proximal algorithms for nonsmooth fractional programs, if $g$ is smooth and $h_{2}\equiv 0$ (i.e., $h=h_{1}$) in problem (1). In this case, we know that $y^{k}=\nabla g(x^{k})$, $z^{k}=0$ and $g(x^{k})\nabla h(x^{k})=\nabla(hg)(x^{k})-h(x^{k})\nabla g(x^{k})$. Using these facts and $c_{k}=1/g(x^{k})$ together with the equivalence between (33) and (36), we can express (33) as
(55) $x^{k+1}\in\arg\min\left\{f(x)+\Big{\langle}\nabla(hg)(x^{k})-\left(\frac{f(x^{k})}{g(x^{k})}+h(x^{k})\right)\nabla g(x^{k}),\right.\\ \left.x-x^{k}\Big{\rangle}+\frac{g(x^{k})}{2\alpha}\|x-x^{k}\|_{2}^{2}:x\in\mathcal{C}\right\}.$
By further choosing $\alpha=g(x^{k})\eta_{k}$ with $\eta_{k}>0$ in (55), we immediately obtain the update formula (6).
5 Sequential convergence of AMPDA
In this section, we investigate the convergence of the entire sequence generated by the AMPDA for solving problem (1) under the KL assumption. To this end, we first present a key proposition regarding the sufficient descent of $Q$ defined by (31).
Proposition 5.1.
Suppose that 2 holds, and let $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ be generated by AMPDA. Then, there exists $a>0$, such that, for all $k\in\mathbb{N}$,
(56) $Q(x^{k+1},y^{k},z^{k},c_{k+1})+a\|x^{k+1}-x^{k}\|_{2}^{2}\leq Q(x^{k},y^{k-1},z^{k-1},c_{k}).$
Proof 5.2.
Thanks to Proposition 4.5, we know that (34) holds with the $\widehat{x}^{k}=x^{k+1}$ for any $k\in\mathbb{N}$. This, along with (32), implies that (56) holds with $a=\sigma/2$.
Next we prove that the relative error condition stated in Proposition 2.2 (ii) holds for $Q$ at $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$ generated by the AMPDA.
Proposition 5.3.
Suppose that Assumption 2 holds and let $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ be generated by AMPDA. Then there exists $b>0$ such that, for all $k\in\mathbb{N}$,
(57) $\mathrm{dist}(0,\partial Q(x^{k+1},y^{k},z^{k},c_{k+1}))\leq b\|x^{k+1}-x^{k}\|_{2}.$
Proof 5.4.
For any $(\bar{x},\bar{y},\bar{z})\in(\Omega\cap\mathcal{C})\times\mathrm{dom}(\partial g^{*})\times\mathrm{dom}(h_{2}^{*})$ and $\bar{c}=1/g(\bar{x})$, we have
(58) ^$\widehat{\partial}Q(\bar{x},\bar{y},\bar{z},\bar{c})\\ \supseteq\left[\begin{array}[]{l}\left(2\bar{c}+\bar{c}^{2}(g^{*}(\bar{y})-\langle\bar{x},\bar{y}\rangle)\right)\widehat{\partial}f_{\mathcal{C}}(\bar{x})-\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\bar{y}+\nabla h_{1}(\bar{x})-\bar{z}\\ \bar{c}^{2}f_{\mathcal{C}}(\bar{x})\partial g^{*}(\bar{y})-\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\bar{x}\\ \widehat{\partial}h_{2}^{*}(\bar{z})-\bar{x}\\ 2f_{\mathcal{C}}(\bar{x})+2\bar{c}f_{\mathcal{C}}(\bar{x})\left(g^{*}(\bar{y})-\langle\bar{x},\bar{y}\rangle\right)\end{array}\right].$
The proof of (58) is given in Appendix A. It is known from the proximal step (33) of AMPDA and the generalized Fermat’s Rule that
(59) $w_{f_{\mathcal{C}}}^{k+1}:=\frac{g(x^{k})}{\alpha_{k}}(x^{k}-x^{k+1}-\alpha_{k}\nabla h_{1}(x^{k})+\alpha_{k}z^{k})+c_{k}f(x^{k})y^{k}\in\widehat{\partial}f_{\mathcal{C}}(x^{k+1}).$
Then, with the help of (58) and the fact that $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$, we have
(60) $w^{k+1}:=(w_{x}^{k+1},w_{y}^{k+1},x^{k}-x^{k+1},w_{c}^{k+1})\in\widehat{\partial}Q\left(x^{k+1},y^{k},z^{k},c_{k+1}\right),$
where $w_{x}^{k+1},w_{y}^{k+1}$ and $w_{c}^{k+1}$ is given as below:
(61) $w_{x}^{k+1}$ $:=\left(2c_{k+1}+c_{k+1}^{2}(g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle)\right)w_{f_{\mathcal{C}}}^{k+1}$
$\qquad-c_{k+1}^{2}f_{\mathcal{C}}(x^{k+1})y^{k}+\nabla h_{1}(x^{k+1})-z^{k},$
(62) $w_{y}^{k+1}$ $:=c_{k+1}^{2}f_{\mathcal{C}}(x^{k+1})(x^{k}-x^{k+1}),$
(63) $w_{c}^{k+1}$ $:=2c_{k+1}f_{\mathcal{C}}(x^{k+1})\left(g(x^{k+1})+g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle\right).$
We then show that $\|w^{k+1}\|_{2}\leq b\|x^{k+1}-x^{k}\|_{2}$ for some $b>0$. It amounts to prove that $w_{x}^{k+1}$, $w_{y}^{k+1}$ and $w_{c}^{k+1}$ can be bounded by the term $\|x^{k+1}-x^{k}\|_{2}$. By invoking Lemma 4.3 (iii), we derive from (62) that
(64) $\|w_{y}^{k+1}\|_{2}\leq M_{f/g^{2}}\|x^{k+1}-x^{k}\|_{2}.$
for all $k\in\mathbb{N}$. By substituting $g^{*}(y^{k})=\langle x^{k},y^{k}\rangle-g(x^{k})$ into (63), we obtain $w_{c}^{k+1}=2c_{k+1}f_{\mathcal{C}}(x^{k+1})(g(x^{k+1})-g(x^{k})+\langle x^{k}-x^{k+1},y^{k}\rangle).$ Invoking this and Lemma 4.3, we deduce that
(65) $\left\|w_{c}^{k+1}\right\|_{2}\leq 2M_{f/g^{2}}M_{g}(L_{g}+M_{\partial g})\|x^{k+1}-x^{k}\|_{2},$
for all $k\in\mathbb{N}$, where $L_{g}>0$ denotes the Lipschitz modulus of $g$ on $\mathcal{X}$. We next focus on the term $w_{x}^{k+1}$. We observe that $w_{x}^{k+1}$ can be rewritten as
(66) $w_{x}^{k+1}=\left(2c_{k+1}+c_{k+1}^{2}(g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle)-c_{k}\right)w_{f_{\mathcal{C}}}^{k+1}\\ +\left(c_{k}^{2}f_{\mathcal{C}}(x^{k})-c_{k+1}^{2}f_{\mathcal{C}}(x^{k+1})\right)y^{k}+\nabla h_{1}(x^{k+1})-\nabla h_{1}(x^{k})+\frac{1}{\alpha_{k}}(x^{k}-x^{k+1}).$
For the term $2c_{k+1}+c_{k+1}^{2}\left(g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle\right)-c_{k}$ in (66), we have
(67) $\left|2c_{k+1}+c_{k+1}^{2}\left(g^{*}(y^{k})-\left\langle x^{k+1},y^{k}\right\rangle\right)-c_{k}\right|$
$=\left|c_{k+1}^{2}c_{k}\left(2g(x^{k+1})g(x^{k})-g^{2}(x^{k})-g^{2}(x^{k+1})\right)+\left\langle x^{k}-x^{k+1},c_{k+1}^{2}y^{k}\right\rangle\right|$
$\leq\left|c_{k+1}^{2}c_{k}\left(g(x^{k+1})-g(x^{k})\right)^{2}\right|+\left\|c_{k+1}^{2}y^{k}\right\|_{2}\|x^{k+1}-x^{k}\|_{2}$
$\leq m_{g}^{-2}\left(2L_{g}+M_{\partial g}\right)\|x^{k+1}-x^{k}\|_{2},$
where the first equality follows from $g^{*}(y^{k})=\langle x^{k},y^{k}\rangle-g(x^{k})$. Besides, in view of boundedness of $\left\{\left(x^{k},y^{k},z^{k},c_{k}\right):k\in\mathbb{N}\right\}$ (see Corollary 4.7) and $\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma\leq\alpha_{k}\leq\overline{\alpha}$ (see Proposition 4.5), we deduce from (59) that
$M_{w_{f_{\mathcal{C}}}}:=\sup\{\|w_{f_{\mathcal{C}}}^{k+1}\|_{2}:k\in\mathbb{N}\}<+\infty.$
Invoking Lemma 4.3, we derive from (66) and (67) that
(68) $\|w_{x}^{k+1}\|_{2}$ $\leq M_{w_{f_{\mathcal{C}}}}m_{g}^{-2}(2L_{g}+M_{\partial g})\|x^{k+1}-x^{k}\|_{2}$
$\quad+\left(L_{f/g^{2}}M_{\partial g}+L_{\nabla h_{1}}+(\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma)^{-1}\right)\|x^{k+1}-x^{k}\|_{2},$
for each $k\in\mathbb{N}$, where $L_{f/g^{2}}>0$ and $L_{\nabla h_{1}}>0$ are the Lipschitz moduli of $f/g^{2}$ and $\nabla h_{1}$ on $\mathcal{X}$, respectively. In view of (60), (64), (65) and (68), we conclude that (57) holds for each $k\in\mathbb{N}$. This completes the proof.
Now we are ready to establish the sequential convergence of the AMPDA.
Theorem 5.5.
Suppose that 2 holds. If the $Q$ given by (31) is a proper closed KL function, then the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA converges to a critical point of (1).
Proof 5.6.
Let the sequence $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ be generated by AMPDA. Then, according to Proposition 2.2, Corollary 4.7, Theorem 4.9, Proposition 5.1 and Proposition 5.3, it remains to prove that $Q_{\infty}:=\lim_{k\to\infty}Q(x^{k+1},y^{k},z^{k},c_{k+1})$ exists and $Q\equiv Q_{\infty}$ holds on the set of all accumulation points of $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$. In view of Step 2 of AMPDA and (32), we have $F(x^{k+1})\leq Q(x^{k+1},y^{k},z^{k},c_{k+1})\leq F(x^{k}).$ This together with Corollary 4.7 leads to the existence of $Q_{\infty}$ and $Q_{\infty}=F_{\infty}$.
We let $(x^{\star},y^{\star},z^{\star},c_{\star})$ be an accumulation point of $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$, and $\{(x^{k_{j}+1},y^{k_{j}},z^{k_{j}},c_{k_{j}+1}):j\in\mathbb{N}\}$ be the subsequence converging to $(x^{\star},y^{\star},z^{\star},c_{\star})$. Firstly, we have $c_{\star}=\lim_{j\to\infty}c_{k_{j}+1}=\lim_{j\to\infty}1/g(x^{k_{j}+1})=1/g(x^{\star})$ due to the continuity of $1/g$ on $\mathcal{X}$. Secondly, we know that $x^{k_{j}}\to x^{\star}$ thanks to $x^{k_{j}+1}\to x^{\star}$ and (51). Thirdly, $(y^{\star},z^{\star})\in\partial g(x^{\star})\times\partial h_{2}(x^{\star})$ follows from $(y^{k_{j}},z^{k_{j}})\to(y^{\star},z^{\star})$ with $(y^{k_{j}},z^{k_{j}})\in\partial g(x^{k_{j}})\times\partial h_{2}(x^{k_{j}})$ for each $j\in\mathbb{N}$. Consequently, we have
$Q(x^{\star},y^{\star},z^{\star},c_{\star})\overset{{(I)}}{=}F(x^{\star})\overset{{(II)}}{=}\lim_{j\to\infty}F(x^{k_{j}+1})\overset{{(III)}}{=}F_{\infty},$
where ${(I)}$ follows from (32), ${(II)}$ holds since $F$ is continuous on the compact $\mathcal{X}\subseteq\Omega$, and ${(III)}$ holds thanks to Corollary 4.7. This completes the proof.
Next, we show that problems (2) and (3) meet the requirements on $Q$, and hence the solution sequence generated by AMPDA for the two problems converges to a critical point of them if the initial point $x^{0}$ satisfies (41). First, for problem (2) $g^{*}$ is the indicator function on the $\ell_{2}$ unit ball, while for problem (3) $g^{*}$ is the indicator function on $\{x\in\mathbb{R}^{n}:\|x\|_{\infty}\leq 1,\|x\|_{1}\leq K\}$ (see, for example, [41]). Second, for problems (2) and (3), the corresponding auxiliary objectives $Q$ (see (31)) are in the form of
(69) $Q(x,y,z,c)=\iota_{\mathcal{C}}(x)+2c\|x\|_{1}+g^{*}(y)\\ -c^{2}\|x\|_{1}\langle x,y\rangle+\frac{\lambda}{2}\|Ax-b\|_{2}^{2}+h_{2}^{*}(z)-\langle x,z\rangle,$
where $\iota_{\mathcal{C}}$ is the indicator function on $\{x\in\mathbb{R}^{n}:\underline{x}\leq x\leq\overline{x}\}$, $g^{*}$ is as discussed above, and $h_{2}(x)=\frac{\lambda}{2}\left\|\mathcal{T}_{\mu}(Ax-b)\right\|_{2}^{2}$. Since $h_{2}$ is convex piecewise linear quadratic, $h_{2}^{*}$ is also piecewise linear quadratic ([34, Theorem 11.14]). Invoking this and the equation (69) together with the definition of semi-algebraic functions ([12, Definition 5]), we know that $Q$ is closed and semi-algebraic, and hence $Q$ is a closed KL function ([12, Theorem 3]). In view of the above display and the discussions right after Proposition 4.1, we finally deduce from Theorem 5.5 the following theorem.
Theorem 5.7.
Consider problems (2) and (3). Then the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA for each of them converges to a critical point of the respective problem if the initial point $x^{0}$ satisfies (41).
6 Numerical experiments
In this section, we conduct some preliminary numerical experiments to evaluate the performance of our proposed AMPDA. All the experiments are conducted in MATLAB R2022a on a desktop equipped with an Intel Core i5-9500 CPU (3.00 GHz) and 16GB of RAM.
In our experiments, we compare the performance of the proposed AMPDA with the gradient descent flow algorithm (GDFA) [38, Algorithm 1] for solving problems (2) and (3). The parameters of AMPDA are set as follows. We set $\underline{\alpha}=10^{-4}$, $\overline{\alpha}=10^{4}$, $\sigma=10^{-5}$, and $\gamma=1/2$ throughout the experiments. Motivated by the Barzilai-Borwein spectral method [3], we update the $\widetilde{\alpha}_{k}$ in Step 1 of AMPDA as
(70) $\widetilde{\alpha}_{k}:=\max\left\{\underline{\alpha},\min\left\{\overline{\alpha},\frac{\|\Delta x^{k}\|_{2}^{2}}{|\langle\Delta x^{k},\Delta h_{1}^{k}\rangle|}\right\}\right\},$
where $\Delta x^{k}=x^{k}-x^{k-1}$ and $\Delta h_{1}^{k}=\nabla h_{1}(x^{k})-\nabla h_{1}(x^{k-1})$ if $k\geq 1$ and $|\langle\Delta x^{k},\Delta h_{1}^{k}\rangle|\neq 0$; otherwise, we set $\widetilde{\alpha}_{k}=1$. Let $T\in\mathbb{R}^{m\times m}$ be the diagonal matrix of which the $j$-th ($j=1,2,...,m$) diagonal entry is set as $T_{j,j}=1$ if $(\mathcal{T}_{\mu}(b))_{j}\neq 0$, and $T_{j,j}=0$ otherwise. For the AMPDA, we pick $i_{\star}\in\arg\max\{|(b-Tb)^{\top}a_{i}|:i\in\{1,2,\cdots,n\}\}$ with $a_{i}$ denoting the $i$-th column of $A$, compute
(71) $\hat{\theta}:=\frac{(b-Tb)^{\top}a_{i_{\star}}}{\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}},$
and then choose the initial point $x^{0}$ as
(72) $x_{i}^{0}=\begin{cases}\arg\min\left\{|\theta-\hat{\theta}|:\theta\in[\underline{x}_{i_{\star}},\overline{x}_{i_{\star}}]\right\},&\text{ if }i=i_{\star},\\ 0,&\text{ else. }\end{cases}$
We show in Appendix B that the $x^{0}$ given by (72) satisfies the condition (41) for both problem (2) and (3), if $A\in\mathbb{R}^{m\times n}~{}(m<n)$ has full row-rank, $b\notin\mathcal{S}_{\mu}$, and $\underline{x}<0<\overline{x}$. For the GDFA, the parameters are set as $\rho=2$, $\beta=10^{-6}$, and $\mathrm{jMax}=5$, following the notations in [38, Algorithm 1]. Also, the GDFA is initialized at the $x^{0}$ given by (72). All the tested algorithms are terminated when
$\frac{\|x^{k}-x^{k-1}\|_{2}}{\|x^{k}\|_{2}}<10^{-6}.$
Following similar settings to [43, Section 7.1], we generate a sensing matrix $A\in\mathbb{R}^{m\times n}$ with i.i.d. standard Gaussian entries and then normalize each column of $A$. Next, a subset of size $K^{\dagger}$ is chosen uniformly at random from $\{1,2,\cdots,n\}$ and a $K^{\dagger}$-sparse vector $x^{\dagger}\in\mathbb{R}^{n}$ with i.i.d. standard Gaussian entries on this subset is generated as the original true signal. The lower bound $\underline{x}$ and the upper bound $\overline{x}$ are set to be $-\max\{5,\|x^{\dagger}\|_{\infty}\}\mathbf{1}_{n}$ and $\max\{5,\|x^{\dagger}\|_{\infty}\}\mathbf{1}_{n}$, respectively, where $\mathbf{1}_{n}$ denotes the $n$-dimensional vector with all entries being $1$. We also produce $\widetilde{z}\in\mathbb{R}^{m}$ as a $\mu^{\dagger}$-sparse vector with $\mu^{\dagger}$ i.i.d. standard Gaussian entries at random (uniformly chosen) positions and set $z=2\mathrm{sign}(\widetilde{z})$. Finally, the measurement $b\in\mathbb{R}^{m}$ is generated as $b=Ax^{\dagger}-z+0.01\epsilon$, where $\epsilon\in\mathbb{R}^{m}$ has i.i.d. standard Gaussian entries.
In our numerical tests, we set $(m,n,K^{\dagger},\mu^{\dagger})=(1280R,365R,40R,5R)$ with $R\in\{1,2,3,4\}$. For each $R$, we generate 50 random instances as described above. The model parameter $\lambda$ is set to $5$ and $0.5$ for problems (2) and (3), respectively. We set $K=\lceil 1.3K^{\dagger}\rceil$ for problem (3), and $\mu=\lceil 1.3\mu^{\dagger}\rceil$ for both problems.
Table 1: Results for solving problem (2).
Algorithm Time Iter Obj RecErr
$R=1$ AMPDA 0.005 39 5.057 2.77E-02
GDFA 0.278 32 5.057 2.77E-02
$R=2$ AMPDA 0.049 49 7.255 2.28E-02
GDFA 1.156 36 7.255 2.28E-02
$R=3$ AMPDA 0.156 55 8.892 2.20E-02
GDFA 2.761 40 8.892 2.18E-02
$R=4$ AMPDA 0.322 64 10.332 2.11E-02
GDFA 5.250 45 10.332 2.10E-02
Table 2: Results for solving problem (3).
Algorithm Time Iter Obj RecErr
$R=1$ AMPDA 0.005 31 1.105 3.68E-02
GDFA 1.068 142 1.105 3.65E-02
$R=2$ AMPDA 0.041 42 1.014 1.72E-02
GDFA 5.307 180 1.014 1.61E-02
$R=3$ AMPDA 0.142 53 1.019 1.98E-02
GDFA 16.098 246 1.019 2.01E-02
$R=4$ AMPDA 0.363 75 1.023 2.09E-02
GDFA 35.859 321 1.023 2.13E-02
The computational results are summarized in Table 1 and Table 2, where we report the computational time in seconds (Time), the number of iterations (Iter), the objective value (Obj), and the recovery error ($\text{RecErr}=\|\widehat{x}-x^{\dagger}\|_{2}/\|x^{\dagger}\|_{2}$ with $\widehat{x}$ being the solution returned by the algorithms) averaged over 50 random instances. One can observe that the objective values and recovery error obtained by the AMPDA are comparable to those of the GDFA, while the AMPDA significantly outperforms the GDFA in terms of computational time for all the problem sizes $R\in\{1,2,3,4\}$. These results demonstrate the efficiency of the proposed AMPDA for solving the scale-invariant sparse signal recovery models (2) and (3).
Appendix A The proof of (58)
In order to prove (58), we first show the following proposition, which concerns the generalized product rule of the Fréchet subdifferential.
Proposition A.1.
Let $\varphi_{l},\varphi_{c},\varphi_{r}:\mathbb{R}^{n}\to(-\infty,+\infty]$, and define $\Phi:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ as follows:
$\Phi(x)=\begin{cases}\varphi_{l}(x)(\varphi_{c}(x)+\varphi_{r}(x)),&\text{ if }x\in\mathrm{dom}(\varphi_{l})\cap\mathrm{dom}(\varphi_{c})\cap\mathrm{dom}(\varphi_{r}),\\ +\infty,&\text{else.}\end{cases}$
Let $\bar{x}\in\mathrm{dom}(\varphi_{l})\cap\mathrm{dom}(\varphi_{c})\cap\mathrm{dom}(\varphi_{r})$, and suppose the following three conditions hold:
  1. (i)
    The $\varphi_{l}$ is nonnegative and locally Lipschitz continuous around $\bar{x}$ relative to $\mathrm{dom}(\varphi_{l})$;
  2. (ii)
    The $\varphi_{c}$ is convex and $\partial\varphi_{c}(\bar{x})\neq\emptyset$;
  3. (iii)
    The $\varphi_{r}$ is continuous at $\bar{x}$ relative to $\mathrm{dom}(\varphi_{r})$.
Then, it holds that
^$\widehat{\partial}\Phi(\bar{x})\supseteq\widehat{\partial}\psi(\bar{x})+\varphi_{l}(\bar{x})\partial\varphi_{c}(\bar{x}),$
where $\psi:\mathbb{R}^{n}\to(-\infty,+\infty]$ is defined as:
$\psi(x)=\begin{cases}(\varphi_{c}+\varphi_{r})(\bar{x})\varphi_{l}(x)+\varphi_{l}(\bar{x})\varphi_{r}(x),&\text{ if }x\in\mathrm{dom}(\varphi_{l})\cap\mathrm{dom}(\varphi_{r}),\\ +\infty,&\text{else.}\end{cases}$
Proof A.2.
From the definition of the Fréchet subdifferential of $\Phi$ at $\bar{x}$, we have
(73) ^$\widehat{\partial}\Phi(\bar{x})=\left\{v:\mathop{\lim\inf}_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\Phi)\end{subarray}}\frac{\Phi(x)-\Phi(\bar{x})-\langle v,x-\bar{x}\rangle}{\|x-\bar{x}\|_{2}}\geq 0\right\}.$
Let $\bar{y}\in\partial\varphi_{c}(\bar{x})$. Due to $\varphi_{l}\geq 0$ around $\bar{x}$, it holds for all $x\in\mathrm{dom}(\Phi)$ that
(74) $\Phi(x)-\Phi(\bar{x})-\langle v,x-\bar{x}\rangle$
$\geq\varphi_{l}(x)(\varphi_{c}(\bar{x})+\langle\bar{y},x-\bar{x}\rangle+\varphi_{r}(x))-(\varphi_{l}(\varphi_{c}+\varphi_{r}))(\bar{x})-\langle v,x-\bar{x}\rangle$
$=\psi(x)-\psi(\bar{x})-\langle v-\varphi_{l}(x)\bar{y},x-\bar{x}\rangle+r(x),$
where $r(x)=(\varphi_{l}(x)-\varphi_{l}(\bar{x}))(\varphi_{r}(x)-\varphi_{r}(\bar{x}))$. It follows that
(75) $\lim_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\Phi)\end{subarray}}\frac{|r(x)|}{\|x-\bar{x}\|_{2}}=0,$
owing to the Lipschitz continuity of $\varphi_{l}$ around $\bar{x}$ relative to $\mathrm{dom}(\varphi_{c})$, and the continuity of $\varphi_{r}$ at $\bar{x}$ relative to $\mathrm{dom}(\varphi_{r})$. Combining (73), (74), and (75), we have
(76) ^$\widehat{\partial}\Phi(\bar{x})$$\supseteq\left\{v:\mathop{\lim\inf}_{\begin{subarray}{c}x\to\bar{x}\\ x\neq\bar{x}\\ x\in\mathrm{dom}(\varphi_{l})\cap\mathrm{dom}(\varphi_{r})\end{subarray}}\frac{\psi(x)-\psi(\bar{x})-\langle v-\varphi_{l}(x)\bar{y},x-\bar{x}\rangle}{\|x-\bar{x}\|_{2}}\geq 0\right\}$
=^$=\widehat{\partial}\psi(\bar{x})+\varphi_{l}(\bar{x})\bar{y},$
for each $\bar{y}\in\partial\varphi_{c}(\bar{x})$. Consequently, this proposition is derived from (76).
Now, we are ready to prove the relation (58). Recall that $(\bar{x},\bar{y},\bar{z})\in(\Omega\cap\mathcal{C})\times\mathrm{dom}(\partial g^{*})\times\mathrm{dom}(h_{2}^{*})$, and $\bar{c}=1/g(\bar{x})>0$ in (58). According to (31), $Q$ around $(\bar{x},\bar{y},\bar{z},\bar{c})$ can be represented as
(77) $Q(x,y,z,c)=c^{2}f_{\mathcal{C}}(x)\Big{(}g^{*}(y)+2/c-\langle x,y\rangle\Big{)}+h_{1}(x)+h_{2}^{*}(z)-\langle x,z\rangle,$
if $(x,y,z,c)\in\mathcal{C}\times\mathrm{dom}(g^{*})\times\mathrm{dom}(h_{2}^{*})\times(0,+\infty)$, and $Q(x,y,z,c)=+\infty$ otherwise. We define $q:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\to(-\infty,+\infty]$ as
(78) $q(x,y,z,c):=c^{2}f_{\mathcal{C}}(x)\left(g^{*}(y)+2/c-\langle x,y\rangle\right)$
if $(x,y,c)\in\mathcal{C}\times\mathrm{dom}(g^{*})\times(0,+\infty)$, and $q(x,y,z,c):=+\infty$ otherwise. Then, we derive from (77) that
(79) ^$\widehat{\partial}Q(\bar{x},\bar{y},\bar{z},\bar{c})=\widehat{\partial}q(\bar{x},\bar{y},\bar{z},\bar{c})+\left[\begin{array}[]{c}\nabla h_{1}(\bar{x})-\bar{z}\\ 0_{n}\\ \widehat{\partial}h_{2}^{*}(\bar{z})-\bar{x}\\ 0\end{array}\right],$
since $h_{1}(x)-\langle x,z\rangle$ is differentiable and $h_{2}$ is separable from $q$. Besides, with the help of Proposition A.1, it follows from (78) that
(80) ^$\widehat{\partial}q(\bar{x},\bar{y},\bar{z},\bar{c})\supseteq\widehat{\partial}q_{2}(\bar{x},\bar{y},\bar{z},\bar{c})+\left[\begin{array}[]{c}0_{n}\\ \bar{c}^{2}f_{\mathcal{C}}(\bar{x})\partial g^{*}(\bar{y})\\ 0_{n}\\ 0\end{array}\right],$
where $q_{2}$ is given by
(81) $q_{2}(x,y,z,c):=(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)c^{2}f_{\mathcal{C}}(x)+\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\left(2/c-\langle x,y\rangle\right),$
if $(x,c)\in\mathcal{C}\times(0,+\infty)$, and $q_{2}(x,y,z,c):=+\infty$ otherwise. On the right-hand side of (81), note that $g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle=g(\bar{x})>0$. Hence, by utilizing Proposition A.1 again and noting that $2/c-\langle x,y\rangle$ is differentiable at $(\bar{x},\bar{y},\bar{c})$, (81) implies that
(82) ^$\widehat{\partial}q_{2}(\bar{x},\bar{y},\bar{z},\bar{c})\supseteq\left[\begin{array}[]{c}(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)\bar{c}^{2}\widehat{\partial}f_{\mathcal{C}}(\bar{x})\\ 0_{n}\\ 0_{n}\\ 2(g^{*}(\bar{y})+2/\bar{c}-\langle\bar{x},\bar{y}\rangle)f_{\mathcal{C}}(\bar{x})\bar{c}\end{array}\right]+\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\left[\begin{array}[]{c}-\bar{y}\\ -\bar{x}\\ 0_{n}\\ -2/\bar{c}^{2}\end{array}\right].$
Combining (79), (80), and (82), we conclude that (58) holds.
Appendix B The initial point $x^{0}$ given by (72) satisfies the condition (41)
Firstly, we shall show $\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}\neq 0$, ensuring that the $\hat{\theta}$ given by (71) is well-defined. Note that $(b-Tb)^{\top}A\neq 0$, since $A$ has full row-rank and $b\neq Tb$ follows from the fact that $b\notin\mathcal{S}_{\mu}$. Then we derive $|(b-Tb)^{\top}a_{i_{\star}}|>0$ immediately, which along with the fact $b^{\top}(a_{i_{\star}}-Ta_{i_{\star}})=(b-Tb)^{\top}a_{i_{\star}}$ implies that $a_{i_{\star}}-Ta_{i_{\star}}\neq 0$. Hence, we deduce that $\|a_{i_{\star}}\|^{2}_{2}-\|Ta_{i_{\star}}\|^{2}_{2}>0$ and $\hat{\theta}\neq 0$. Secondly, due to $\underline{x}<0<\overline{x}$ and $\hat{\theta}\neq 0$, the $x^{0}$ given by (72) is feasible for problems (2) and (3), that is, $x^{0}\neq 0$ and $\underline{x}\leq x^{0}\leq\overline{x}$ hold. Thirdly, owing to $0\in[\underline{x}_{i_{\star}},\overline{x}_{i_{\star}}]$ and $x_{i_{\star}}^{0}\neq 0$, we deduce from (72) that $(x_{i_{\star}}^{0}-\hat{\theta})^{2}<\hat{\theta}^{2}$, or equally, $(x_{i_{\star}}^{0}-2\hat{\theta})x_{i_{\star}}^{0}<0$. Note that $\|\mathcal{T}_{\mu}(Ax-b)\|^{2}_{2}=\max\left\{\|E(Ax-b)\|^{2}_{2}:E\in\{0,1\}^{m\times m}\text{ is diagonal,}~{}\|E\|_{0}\leq\mu\right\},$ where $\|\cdot\|_{0}$ counting the number of nonzero elements in the matrix. Then, we have
$F(x^{0})\leq\frac{\|x^{0}\|_{1}}{\|x^{0}\|_{2}}+\frac{\lambda}{2}\|Ax^{0}-b\|_{2}^{2}-\frac{\lambda}{2}\|T(Ax^{0}-b)\|_{2}^{2}$
$=1+\frac{\lambda}{2}\left(\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}\right)(x_{i_{\star}}^{0}-2\hat{\theta})x_{i*}^{0}+\frac{\lambda}{2}\left(\|b\|_{2}^{2}-\|Tb\|_{2}^{2}\right)$
<$\overset{(*)}{<}1+\frac{\lambda}{2}(\|b\|_{2}^{2}-\|Tb\|_{2}^{2})=1+\frac{\lambda}{2}\mathrm{dist}^{2}(b,\mathcal{S}_{\mu}),$
where ()(*)( ∗ ) follows from the positivity of $\|a_{i_{\star}}\|_{2}^{2}-\|Ta_{i_{\star}}\|_{2}^{2}$ and the negativity of $(x_{i_{\star}}^{0}-2\hat{\theta})x_{i_{\star}}^{0}$. This means that $x^{0}$, defined in (72), satisfies the condition (41).